### v1

# XAMPLER: Learning to Retrieve Cross-Lingual In-Context Examples

Abstract
Recent studies have shown that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving high-quality in-context examples, significantly improves in-context learning of English. However, adapting these methods to other languages, especially low-resource ones, presents challenges due to the scarcity of available cross-lingual retrievers and annotated data. In this paper, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data. XAMPLER first trains a retriever with positive/negative English samples, which are constructed based on the predictions of the multilingual large language model for in-context learning. Then, the trained retriever is directly employed to retrieve English examples as few-shot examples for in-context learning of target languages. Experiments on the massively multilingual text classification benchmark of SIB200 with 176 languages demonstrate that XAMPLER substantially improves the in-context learning performance across languages. Our code is available at https://github.com/cisnlp/XAMPLER.
1 Introduction
Large language models (LLMs) have shown emergent abilities in in-context learning, where a few input-output examples are provided with the input query. Through in-context learning, LLMs can yield promising results without any parameter updates (Brown et al., 2020). However, the efficacy of in-context learning is highly dependent on the selection of the few-shot examples (Liu et al., 2022).
Recent studies (Luo et al., 2024) have uncovered a more strategic approach to example retrieval. Rather than relying on random selection, these studies advocate for retrieving examples tailored to the input query, resulting in notable performance enhancements in in-context learning. The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals using labeled data.
Utilizing off-the-shelf retrievers has been further validated as an effective approach in multilingual settings (Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023). However, this method encounters limitations when applied to low-resource languages. Existing multilingual retrievers, e.g., (Reimers and Gurevych, 2020), cover a limited number of languages (i.e., 50+), and language-model-based retrievers (Hu et al., 2020) struggle to effectively align distant languages (Cao et al., 2020; Liu et al., 2023). Additionally, relying on off-the-shelf retrievers might lead to sub-optimal performance. Conversely, adopting task-specific fine-tuned retrievers has been demonstrated as a more effective approach (Rubin et al., 2022). Nonetheless, the availability of data for fine-tuning task-specific retrievers in low-resource languages is limited.
To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval). Specifically, given an English query and an English candidate training example as input, we employ in-context learning with MaLA500 (Lin et al., 2024b), a massively multilingual LLM covering 534 languages, to predict the label of the query. Based on the correctness of the prediction, we classify the candidate example as either positive or negative. Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a massively multilingual language model covering 534 languages, with the objective of minimizing the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023). Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language. The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for further in-context learning. Our extensive experiments across a range of 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning performance across languages.
2 Approach
2.1 Problem Definition
Given an input query $q_{i}$ in any language, our aim is to retrieve tailored few-shot examples from the sample pool $D$ to enhance the in-context learning of $q_{i}$. Given the scarcity of annotated data in low-resource languages, we introduce XAMPLER, namely, Cross-Lingual Example Retrieval. On one hand, we leverage in-domain English examples as the sample pool $D$, from which we retrieve cross-lingual examples in English for $q_{i}$ in any target language. On the other hand, we only consider $q_{i}$ sourced from English training data to train the task-specific retriever, and then directly apply the retriever for evaluation across different languages.
2.2 Data Construction
To train the task-specific retriever aimed at retrieving informative samples for the given query $q_{i}$, we consider contrastive learning, which requires both positive and negative examples for each query $q_{i}$. We define examples as positive when the LLM accurately predicts the ground truth of $q_{i}$ while utilizing the candidate as a one-shot example appended to $q_{i}$ for in-context learning. Conversely, examples are categorized as negative if the LLM’s prediction deviates from the ground truth.
Scoring all pairs of training examples presents a quadratic complexity in $|D|$, making it resource-intensive. Inspired by Rubin et al. (2022), we mitigate this by selecting the top $k$ similar examples as candidates. We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] This paper uses the distiluse-base-multilingual-cased-v1 version for SBERT. for candidate selection. Based on our experiments detailed in Section C, we set $k=10$. The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
After obtaining the candidate-query pairs $\{(q_{i},d_{i,1}),\cdots,(q_{i},d_{i,k})\}$, we conduct 1-shot in-context learning with MaLA500 (Lin et al., 2024b) to predict the class of the $q_{i}$ given the candidate $d_{i,j}$, resulting in a predicted label $\hat{y_{i,j}}$. If MaLA500 correctly predicts the label of $q_{i}$ (i.e., $\hat{y_{i,j}}=y_{i}$), we consider the candidate $d_{i,j}$ as a positive sample ($d^{+}_{i,j}$); otherwise, it is treated as a negative sample ($d^{-}_{i,j}$). Finally, we divide $D_{i}$ into sets of positive and negative samples, denoted as $D_{i}^{pos}$ and $D_{i}^{neg}$, respectively.
2.3 Retriever Fine-tuning
We utilize a contrastive loss function (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between positive example pairs while minimizing the similarity between negative example pairs. We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant costs of fine-tuning a LLM.
2.4 In-Context Learning
When employing in-context learning across languages, where $q_{i}$ can be in any language, we utilize the fine-tuned task-specific retriever to retrieve a few cross-lingual examples in English tailored to $q_{i}$. The retrieved examples are then appended to $q_{i}$ as input for MaLA500 (Lin et al., 2024b) to predict the label of $q_{i}$ through in-context learning.
3 Experiment
3.1 Setup
Benchmark
We evaluate XAMPLER on a massively multilingual text classification benchmark, i.e., SIB200 (Adelani et al., 2023). SIB200 involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500. For a detailed language list, refer to §A.
Our evaluation framework follows the prompt template used in Lin et al. (2024b): ‘The topic of the news [sent] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources. MaLA500 takes in the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set. We measure the performance with accuracy.
Baselines
We compare XAMPLER with three types of baselines.
  • Random Sampling. We randomly select examples from the training set $D$.
  • Off-the-shelf Retriever. We utilize SBERT (Reimers and Gurevych, 2019), a model covering 50+ languages trained with parallel corpora based on mBERT (Devlin et al., 2019). Additionally, we employ two massively multilingual language models, namely Glot500 and MaLA500, as retrievers, denoted as Glot500 RET and MaLA500 RET, respectively. Tailored examples are retrieved based on the cosine similarity between the sentence representations of the candidate and the query. For Glot500, we utilize mean pooling over hidden states of the selected layer. For MaLA500, we adopt a position-weighted mean pooling method on the selected layer, assigning higher weights to later tokens (Muennighoff, 2022). We use K-Nearest Neighbors (KNN) to select the layer that performs best across layers (See §B). The selected layers for Glot500 and MaLA500 are 11 and 21, respectively.
  • Cross-lingual Transfer. We delve into another category of baselines called cross-lingual transfer, which also harnesses English data. In this approach, the language model is fine-tuned with English data and subsequently deployed for evaluation across target languages. Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines are denoted as Glot500 XLT and MaLA500 XLT. For Glot500, we opt for full-parameter fine-tuning. For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the parameters of the LoRA modules with prompt tuning.
In addition, we consider K-Nearest Neighbors (KNN) with XAMPLER and the baselines with retrievers for comparison. Specifically, we adopt majority voting based on the labels of the retrieved examples with the given retriever.
KNN ICL
Latin Non-Latin Avg Latin Non-Latin Avg
Random - - - 56.38 58.66 57.14
SBERT 44.90 37.13 42.29 63.90 62.93 63.57
Glot500 RET 51.16 60.26 54.21 65.21 70.09 66.85
MaLA500 RET 31.90 32.63 32.15 61.02 63.58 61.88
Glot500 XLT - - - 67.09 74.30 69.51
MaLA500 XLT - - - 69.15 71.39 69.90
XAMPLER 67.05 75.97 70.04 73.59 79.80 75.67
Table 1: Average macro-accuracy across all 176 languages, 117 languages in the Latin script and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines with K-Nearest Neighbors (KNN) or In-Context Learning (ICL). The provided results are all based on the 3-shot setting.
3.2 Main Results
The comparison between the baselines and XAMPLER is illustrated in Table 1. Our analysis reveals several insights based on the performance with In-Context Learning (ICL) across different methods. Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning. Leveraging an off-the-shelf retriever notably boosts performance. For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever. Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%, respectively.
Among all the methods, XAMPLER achieves the highest performance, surpassing the second-best method, MaLA500 XLT, by 5.77%. This highlights the effectiveness of training a task-specific retriever solely with English data. Moreover, XAMPLER outperforms MaLA500 XLT by 4.44% on languages written in Latin scripts and by 8.41% on languages written in non-Latin scripts. It shows that XAMPLER can provide greater benefits for languages with non-Latin scripts, which are relatively isolated from English written in Latin.
3.3 Effect of Task-Specific Retriever
The results from various retrievers utilizing KNN, as shown in Table 1, indicate that XAMPLER’s fine-tuned retriever excels at retrieving more examples within the same classes as the query in the target language. Notably, XAMPLER with KNN outperforms the second-best retriever, Glot500 RET, by a notable margin, i.e., 15.83%. This superiority enables XAMPLER to leverage majority label bias in in-context learning (Zhao et al., 2021), thereby enhancing overall performance.
3.4 Effect of In-Context Learning
In Table 1, we compare the performance of XAMPLER with KNN against XAMPLER with ICL. Notably, XAMPLER with ICL exhibits a superior performance, outperforming KNN by 5.63%. Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages written in non-Latin scripts. This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
We further compare the performance of XAMPLER with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 1. Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%. Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL reaches its peak with only 2 examples. Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%. These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
Figure 1: KNN (K-Nearest Neighbors) vs. ICL (In-Context Learning) with different number of shots. X-axis: the number of retrieved examples. Y-axis: Macro-average accuracy.
4 Related Work
Early studies (Gao et al., 2021; Liu et al., 2022; Rubin et al., 2022) on retrieving informative examples for few-shot in-context learning often rely on off-the-shelf retrievers to gather semantically similar examples to the query.
While off-the-shelf retrievers have shown promise, the examples they retrieve may not always represent optimal solutions for the given task, potentially resulting in sub-optimal performance. Hence, Rubin et al. (2022) delves into learning-based approaches: if an LLM finds an example useful, the retriever should be encouraged to retrieve it. This approach enables direct training of the retriever using signals derived from query and example pairs in the task of interest.
Retrieval-based in-context learning has demonstrated effectiveness in natural language understanding (Liu et al., 2022; Ye et al., 2023), reasoning (Zhang et al., 2023), knowledge-based question answering (Liu et al., 2022; Ye et al., 2023), and text generation (Rubin et al., 2022).
A study closely related to ours is conducted by Shi et al. (2022), which trains a cross-lingual exemplar retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task. However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al. (2022).
5 Conclusion
In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language. By relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language. Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.
Limitation
In this paper, we utilize Glot500 as the base model for training XAMPLER’s retriever, MaLA500 as the LLM for in-context learning, and SIB200 as the benchmark to assess the efficacy of XAMPLER across multiple languages. We opted not to consider alternative models and benchmarks due to the absence of massively multilingual models and benchmarks.
Additionally, while it is acknowledged that English may not universally serve as the optimal source language for cross-lingual transfer across all target languages (Lin et al., 2019; Wang et al., 2023; Lin et al., 2024a), our study does not explore the selection of different source languages due to the predominant availability of training data in English for many tasks.
Appendix A Detailed Results
The language list of SIB200 and the results of XAMPLER and the compared baselines are shown in Table 2 and Table 3.
Random SBERT (KNN) SBERT Glot500 RET (KNN) Glot500 RET MaLA500 RET (KNN) MaLA500 RET Glot500 XLT MaLA500 XLT XAMPLER (KNN) XAMPLER
ace_Latn 61.76 40.20 64.71 50.00 69.12 28.92 69.12 65.12 66.18 75.49 75.98
acm_Arab 55.39 79.90 79.41 64.22 73.04 35.29 62.75 76.81 74.51 76.47 79.90
afr_Latn 65.20 75.00 83.33 60.78 78.92 43.63 72.06 76.10 78.92 82.84 86.27
ajp_Arab 58.82 76.47 79.41 62.25 71.57 32.84 63.24 75.44 72.55 82.84 83.33
als_Latn 64.71 46.57 75.49 63.73 73.53 36.76 67.65 76.76 81.86 65.20 83.33
amh_Ethi 53.92 29.41 58.33 46.57 55.39 22.55 57.35 68.46 61.27 71.57 73.53
apc_Arab 59.80 79.41 79.41 62.25 74.51 31.37 62.25 77.52 75.00 82.84 84.80
arb_Arab 58.33 83.33 81.37 61.76 70.10 36.76 67.16 77.40 78.43 78.43 85.78
ary_Arab 56.37 79.41 78.92 55.39 67.65 25.98 56.37 74.90 67.16 80.39 80.88
arz_Arab 55.39 74.51 74.02 62.25 69.61 33.33 63.73 77.08 71.08 78.92 80.88
asm_Beng 61.27 22.55 61.76 59.80 77.45 22.55 66.67 74.93 76.96 78.92 84.80
ast_Latn 70.59 73.04 81.86 63.73 79.90 49.51 81.37 80.81 84.31 75.49 89.71
ayr_Latn 37.25 24.51 40.69 32.84 41.67 19.12 42.65 49.93 48.53 38.24 50.98
azb_Arab 44.61 26.47 47.55 41.18 59.31 23.04 47.55 65.56 58.82 72.55 73.04
azj_Latn 66.18 58.33 78.92 72.55 76.47 20.10 68.63 78.73 87.25 75.49 85.78
bak_Cyrl 58.33 50.98 74.02 59.80 72.55 28.92 63.73 77.87 75.00 72.55 80.88
bam_Latn 36.76 31.37 42.65 31.86 43.14 19.61 41.18 49.71 44.12 46.08 46.08
ban_Latn 66.18 44.61 70.59 53.92 70.59 31.86 72.06 73.01 72.06 75.98 82.35
bel_Cyrl 64.22 38.24 72.06 59.31 76.96 43.63 73.04 77.16 79.90 69.12 84.31
bem_Latn 42.16 28.43 43.63 32.35 50.00 18.14 39.71 56.89 58.33 60.78 61.76
ben_Beng 63.73 20.10 61.27 60.29 75.00 32.35 69.12 73.95 80.88 79.90 82.35
bjn_Latn 63.73 31.37 62.75 59.31 70.10 33.33 70.59 69.80 69.61 76.96 81.86
bod_Tibt 44.61 09.31 33.33 38.24 48.53 19.61 47.55 62.79 47.06 54.90 59.31
bos_Latn 70.10 55.39 77.45 72.55 79.90 46.08 75.00 81.99 83.33 71.57 87.75
bul_Cyrl 65.69 75.00 82.84 67.16 80.39 52.45 78.92 78.87 81.86 79.90 85.29
cat_Latn 69.61 74.51 84.31 64.22 74.51 61.27 80.88 77.97 86.76 71.57 88.73
ceb_Latn 66.18 45.59 73.53 63.73 80.88 31.37 72.55 78.80 81.86 82.35 85.29
ces_Latn 69.61 53.92 76.96 63.24 76.47 55.88 78.92 77.25 86.27 76.96 88.24
cjk_Latn 37.75 28.43 42.16 26.96 45.10 21.57 44.61 48.41 53.43 44.61 47.06
ckb_Arab 57.84 17.65 61.27 54.90 72.06 23.53 61.76 74.46 75.00 79.90 81.37
cmn_Hani 68.63 85.29 83.82 75.98 80.88 62.75 78.92 78.55 82.84 87.25 88.73
crh_Latn 58.82 67.65 75.00 54.41 70.59 14.71 58.33 67.84 71.08 64.71 75.98
cym_Latn 63.24 24.02 65.69 49.02 73.04 27.94 70.10 72.50 77.45 72.55 81.37
dan_Latn 68.14 47.06 78.92 67.65 79.90 56.37 78.43 79.68 84.80 86.27 89.71
deu_Latn 70.59 83.82 86.27 70.10 77.45 57.84 80.88 78.48 87.25 80.88 87.75
dyu_Latn 41.67 28.92 45.59 27.94 50.49 24.51 46.57 45.17 46.57 43.14 48.53
dzo_Tibt 31.37 09.31 20.59 47.55 41.18 16.67 32.84 62.21 33.33 68.63 52.94
ell_Grek 67.65 31.37 76.47 59.80 72.55 40.20 72.06 74.71 79.90 75.98 83.82
eng_Latn 70.59 84.31 87.25 79.41 86.27 69.61 81.86 83.65 85.29 89.22 91.18
epo_Latn 65.69 57.84 75.98 61.27 76.96 33.33 68.63 75.20 77.94 78.92 82.84
est_Latn 61.76 44.61 66.67 65.69 75.49 30.88 66.18 74.07 76.47 67.65 81.37
eus_Latn 57.35 47.55 71.08 66.67 76.47 29.90 68.14 76.08 70.10 66.67 82.84
ewe_Latn 38.24 26.47 39.22 27.94 39.71 20.59 38.73 47.65 46.57 50.98 53.92
fao_Latn 56.86 31.86 58.33 45.10 63.24 29.90 60.78 76.15 69.61 82.35 81.86
fij_Latn 41.18 32.35 49.02 31.86 46.08 21.08 42.65 55.74 51.47 50.00 55.39
fin_Latn 65.20 36.76 71.57 62.75 74.02 48.53 74.51 75.98 80.88 76.96 83.82
fon_Latn 39.22 20.10 41.67 33.33 45.59 20.10 42.65 44.41 45.59 41.67 48.53
fra_Latn 69.61 86.27 84.31 60.78 79.41 63.24 79.90 81.20 85.29 85.78 89.22
ful_Latn 37.75 35.29 48.04 23.53 48.04 21.57 45.59 45.54 45.10 47.06 50.49
fur_Latn 60.29 63.73 72.55 54.90 67.16 36.76 70.59 67.75 73.53 75.49 79.90
gla_Latn 54.90 17.16 52.45 39.22 59.80 25.00 61.76 59.00 60.29 59.31 63.73
gle_Latn 57.84 23.53 57.84 49.02 65.20 31.86 62.75 62.94 69.12 70.10 74.02
glg_Latn 72.06 76.96 81.86 67.65 80.88 53.43 79.90 79.36 84.31 81.37 87.75
grn_Latn 53.92 62.75 69.61 43.14 61.76 27.45 60.78 69.85 64.71 76.47 75.00
guj_Gujr 59.80 13.73 53.43 68.14 73.04 32.84 65.20 77.87 73.53 81.86 86.27
hat_Latn 61.76 50.49 72.06 59.31 74.02 25.49 68.14 73.55 81.86 79.41 82.84
hau_Latn 56.86 23.04 55.88 47.06 60.78 23.53 57.35 62.33 70.10 58.82 70.10
heb_Hebr 45.59 26.96 47.06 58.33 58.33 29.90 43.63 72.60 56.86 75.00 75.00
hin_Deva 60.29 12.75 56.86 67.65 71.08 34.80 65.20 76.84 76.47 84.31 85.78
hne_Deva 56.86 15.69 54.41 55.88 70.10 28.43 66.67 70.07 75.49 79.41 80.88
hrv_Latn 70.10 52.94 78.43 73.53 82.35 51.47 74.51 81.86 85.78 71.57 88.24
hun_Latn 64.71 38.24 66.18 69.61 81.86 50.49 72.06 81.20 84.80 83.82 85.78
hye_Armn 67.16 16.18 67.16 60.78 70.59 31.37 68.14 76.62 77.45 80.88 83.33
ibo_Latn 57.35 37.25 63.24 49.02 70.10 23.53 56.86 69.09 72.06 74.02 79.41
ilo_Latn 61.76 50.49 69.12 47.55 68.14 32.35 65.69 70.81 77.45 75.00 79.90
ind_Latn 69.61 64.22 82.84 75.98 81.37 50.98 75.98 80.34 83.82 88.24 89.71
isl_Latn 60.78 31.86 67.65 55.39 69.61 24.51 62.75 73.04 75.49 77.94 78.92
ita_Latn 71.08 82.84 88.24 68.14 81.37 64.22 77.94 78.80 86.76 79.90 88.24
jav_Latn 64.22 40.69 71.08 56.37 75.00 30.88 70.10 73.19 75.00 79.90 82.35
jpn_Jpan 71.08 76.96 87.25 68.14 76.96 65.69 82.84 78.33 80.88 83.33 87.75
kab_Latn 28.43 20.10 27.94 27.45 36.27 22.06 29.41 38.70 27.94 31.37 37.75
kac_Latn 35.29 28.92 41.18 40.20 42.65 20.10 36.76 54.09 45.10 40.69 46.57
kam_Latn 38.73 30.88 47.55 34.31 46.08 17.65 34.31 46.99 50.49 46.08 50.00
kan_Knda 58.82 15.69 56.37 64.22 71.57 24.02 59.80 74.73 73.04 75.00 77.94
kat_Geor 65.20 15.69 58.82 63.73 75.98 27.94 70.10 77.79 75.98 72.55 83.33
kaz_Cyrl 61.76 43.14 68.63 69.12 74.51 31.37 64.71 75.47 78.43 67.16 79.41
kbp_Latn 37.25 25.00 43.14 34.31 44.61 17.65 44.12 49.83 47.06 41.18 48.04
kea_Latn 64.71 61.27 78.92 48.53 75.49 33.82 75.98 66.81 80.39 74.51 79.41
khm_Khmr 69.12 34.31 69.61 58.82 76.47 38.24 74.51 76.42 79.41 82.84 86.76
kik_Latn 46.57 33.33 51.47 43.63 50.98 19.12 44.12 55.47 52.94 58.82 59.31
kin_Latn 45.59 26.96 50.00 39.22 51.47 19.61 48.53 57.01 67.65 64.71 64.22
kir_Cyrl 58.82 42.65 67.65 68.63 75.98 24.51 55.39 75.93 75.00 63.73 77.45
kmb_Latn 36.76 24.51 41.67 28.43 42.65 24.02 42.65 44.80 48.04 46.57 50.98
kmr_Latn 51.96 27.45 58.82 50.00 69.12 20.10 56.37 63.55 66.18 52.94 70.59
kon_Latn 49.02 50.49 63.24 44.61 63.73 20.10 54.41 58.85 61.27 57.84 64.22
kor_Hang 69.61 81.37 85.29 62.25 78.43 48.53 78.43 77.33 79.90 82.35 87.25
lao_Laoo 64.71 33.33 69.12 58.33 74.02 30.39 70.59 77.33 75.00 79.90 84.31
lij_Latn 64.22 61.76 77.45 50.49 69.61 35.29 66.67 70.07 76.96 70.59 80.39
lim_Latn 61.76 68.63 79.41 55.88 72.55 38.24 72.55 69.49 65.20 71.08 75.98
lin_Latn 48.04 46.57 61.27 48.53 61.27 20.59 54.90 64.95 58.33 63.24 68.63
Table 2: 3-shot accuracy with KNN or ICL (Part I). Methods with ‘(KNN)’ denote that KNN is used with retrieved samples, otherwise ICL is used.
Random SBERT (KNN) SBERT Glot500 RET (KNN) Glot500 RET MaLA500 RET (KNN) MaLA500 RET Glot500 XLT MaLA500 XLT XAMPLER (KNN) XAMPLER
lit_Latn 61.27 41.18 64.71 62.25 72.55 29.41 63.24 78.04 79.41 69.61 85.29
lmo_Latn 63.24 57.35 73.04 48.53 66.67 33.33 66.18 66.25 73.53 72.55 77.45
ltz_Latn 62.75 60.78 76.96 56.37 72.55 35.29 70.10 69.26 80.39 72.06 79.41
lua_Latn 38.24 42.65 50.49 36.27 45.10 24.51 47.06 55.51 52.45 47.55 53.43
lug_Latn 38.24 29.41 41.67 33.82 50.00 20.10 40.69 53.28 51.47 52.45 58.33
luo_Latn 37.75 26.47 40.69 31.86 43.63 18.63 44.12 51.03 50.00 55.39 59.31
lus_Latn 48.53 41.18 57.84 39.71 50.98 30.39 51.96 60.83 61.76 65.69 68.14
lvs_Latn 64.22 39.22 68.63 62.25 73.53 31.86 64.22 78.87 79.90 71.08 82.84
mai_Deva 57.35 17.16 50.49 62.25 69.61 28.92 59.31 76.91 68.63 83.82 78.92
mal_Mlym 57.84 15.69 52.45 58.33 67.16 26.47 59.80 73.33 68.63 77.94 78.92
mar_Deva 59.31 15.69 55.88 60.78 75.00 27.94 63.73 75.96 77.94 73.04 79.41
min_Latn 67.65 44.61 66.67 58.82 72.55 32.84 74.02 70.42 78.43 78.43 79.90
mkd_Cyrl 68.63 57.84 77.94 67.16 75.98 49.02 75.49 76.50 80.88 72.55 83.82
mlt_Latn 68.14 50.00 72.55 60.29 77.45 31.86 67.65 77.87 81.37 78.92 84.31
mon_Cyrl 58.82 21.08 59.31 63.73 73.53 25.00 62.25 75.12 71.57 79.41 82.84
mos_Latn 37.75 39.22 46.08 23.04 39.22 21.08 42.65 47.06 45.10 44.61 50.49
mri_Latn 52.94 22.55 54.41 18.63 46.08 17.16 46.08 54.49 60.78 59.80 68.63
mya_Mymr 53.92 16.67 49.02 53.92 66.18 23.04 50.98 73.68 54.90 74.02 80.39
nld_Latn 68.63 82.84 85.29 75.00 81.37 56.37 81.37 78.24 86.27 87.25 88.73
nno_Latn 68.14 46.57 73.53 64.22 75.49 42.65 74.51 78.63 83.82 86.27 89.22
npi_Deva 63.24 15.20 59.31 67.16 74.51 29.41 70.10 77.99 75.49 82.35 87.75
nso_Latn 41.67 30.39 52.45 39.22 50.49 21.57 43.63 60.71 55.39 54.41 58.82
nya_Latn 45.10 35.29 52.94 50.49 61.76 23.04 49.02 68.36 58.82 66.67 68.14
oci_Latn 67.65 69.61 81.86 58.33 70.10 44.12 77.45 76.50 84.31 74.02 85.78
orm_Latn 34.80 13.73 34.80 40.69 42.16 15.69 35.29 50.59 46.08 42.65 48.53
ory_Orya 51.47 29.41 47.06 59.31 67.65 23.53 57.35 73.55 62.25 80.88 78.92
pag_Latn 55.88 63.24 72.06 59.31 71.57 26.47 62.25 73.31 75.98 80.39 80.88
pan_Guru 57.35 14.22 53.92 51.96 67.16 26.47 60.78 70.44 73.04 71.08 75.98
pap_Latn 65.20 63.73 75.49 58.82 76.96 34.31 70.10 72.23 78.43 77.94 80.39
pes_Arab 66.67 37.25 67.65 69.61 76.96 34.31 71.08 77.72 79.41 87.75 86.76
plt_Latn 52.94 29.41 57.35 42.65 60.29 21.08 50.49 64.22 69.61 48.04 67.16
pol_Latn 68.14 78.43 87.25 67.16 78.92 57.84 78.43 77.97 84.80 75.00 85.29
por_Latn 72.06 81.86 87.75 60.78 81.37 59.31 83.33 79.71 85.78 85.78 90.69
prs_Arab 65.20 32.84 64.22 62.75 76.96 31.86 71.57 78.70 78.92 86.76 86.27
pus_Arab 49.51 25.98 52.94 53.43 59.80 23.53 50.98 69.34 62.25 60.78 71.08
quy_Latn 49.02 29.41 52.45 42.65 52.94 19.61 53.92 59.31 57.84 56.37 62.25
ron_Latn 67.65 58.33 77.94 66.18 75.98 55.39 76.96 78.14 83.33 80.88 86.76
run_Latn 43.14 22.06 44.61 41.18 49.51 18.14 45.59 51.86 60.78 62.75 66.18
rus_Cyrl 69.61 83.82 86.27 72.06 79.90 56.86 77.45 79.53 81.37 86.27 89.71
sag_Latn 43.14 39.71 49.02 39.71 54.41 20.59 45.59 52.50 49.02 56.37 59.31
san_Deva 51.96 15.20 52.94 48.04 59.80 23.53 59.31 70.34 60.29 66.67 68.63
scn_Latn 68.14 55.39 77.45 49.51 71.57 34.31 73.53 67.50 74.51 72.06 80.39
sin_Sinh 61.27 22.06 62.25 66.18 73.53 26.47 65.20 74.29 72.55 79.90 82.84
slk_Latn 66.67 49.51 70.59 63.73 74.02 45.10 75.98 77.99 81.86 74.02 84.31
slv_Latn 63.73 49.02 74.51 64.22 74.02 51.96 75.00 75.51 78.92 66.67 81.86
smo_Latn 54.41 32.84 63.24 36.76 62.75 18.63 52.94 69.14 63.73 66.18 73.04
sna_Latn 45.10 24.51 47.06 42.16 51.47 21.57 40.20 59.31 60.29 55.39 62.75
snd_Arab 44.12 29.90 49.51 50.49 52.45 20.10 44.61 65.83 47.55 59.80 67.16
som_Latn 42.65 20.59 42.65 49.02 59.80 17.16 42.16 56.52 63.24 50.00 58.82
sot_Latn 45.10 24.51 49.02 43.63 58.33 21.57 48.04 63.63 57.84 64.71 67.65
spa_Latn 71.57 84.80 85.78 67.16 78.92 59.80 80.39 79.14 85.29 82.84 90.20
srd_Latn 63.73 64.22 75.98 49.51 70.59 34.80 71.08 66.10 66.67 61.76 78.43
srp_Cyrl 68.63 55.88 79.41 65.69 82.84 47.55 79.41 78.53 82.35 67.65 80.88
ssw_Latn 45.10 24.51 44.12 40.69 56.86 26.96 51.96 63.19 59.80 59.80 63.24
sun_Latn 68.14 43.14 70.10 63.24 78.43 29.90 73.53 75.93 81.37 82.84 85.29
swe_Latn 66.18 46.08 72.55 71.57 77.94 59.31 77.94 79.73 85.29 84.80 87.75
swh_Latn 61.27 28.92 60.29 56.86 66.67 12.75 54.90 72.94 76.96 64.22 75.98
szl_Latn 62.75 67.65 75.98 54.41 70.10 40.69 71.57 68.65 75.49 63.73 77.45
tam_Taml 54.90 15.20 50.98 64.22 71.08 27.45 62.25 75.37 75.98 80.88 78.92
tat_Cyrl 61.76 46.57 71.57 63.73 75.00 24.51 68.63 79.17 78.92 70.59 82.35
tel_Telu 57.84 18.63 50.98 65.69 70.59 25.98 60.29 74.53 74.51 84.80 81.37
tgk_Cyrl 60.29 26.96 61.27 63.73 73.53 35.29 63.73 77.08 80.88 82.35 80.39
tgl_Latn 68.14 38.73 69.61 63.24 79.41 30.88 70.10 78.21 83.82 83.33 86.76
tha_Thai 65.20 25.98 56.86 65.20 75.49 41.18 67.65 77.43 70.10 79.90 86.76
tir_Ethi 44.12 29.41 51.47 42.16 48.53 18.63 47.06 55.91 50.00 55.39 57.35
tpi_Latn 65.20 53.92 75.98 57.84 70.59 27.94 64.71 76.94 74.51 81.37 84.80
tsn_Latn 49.02 31.37 49.02 35.78 50.49 20.10 42.16 56.76 56.86 54.90 59.80
tso_Latn 45.59 25.98 44.12 38.73 51.47 22.06 50.98 56.23 55.39 55.88 62.25
tuk_Latn 53.43 50.00 68.63 51.47 69.61 14.22 50.98 73.33 66.67 61.27 77.94
tum_Latn 42.16 26.96 47.06 32.35 56.86 24.02 43.63 62.55 52.94 65.20 64.71
tur_Latn 68.14 82.84 86.27 68.63 76.96 15.69 56.86 78.06 81.37 73.53 84.80
uig_Arab 45.10 12.75 40.20 55.39 53.43 18.63 37.25 72.18 49.51 72.55 69.12
ukr_Cyrl 66.18 59.31 77.45 67.16 79.90 56.86 79.41 78.63 81.86 76.47 86.27
umb_Latn 33.82 29.90 37.75 33.82 46.57 22.06 39.71 48.97 49.02 47.06 49.51
urd_Arab 55.88 29.90 54.90 59.80 69.12 36.27 66.67 74.29 69.12 64.22 75.98
uzb_Latn 57.35 38.24 62.25 59.80 73.04 24.02 62.75 73.14 76.96 60.78 77.45
vec_Latn 67.65 68.14 82.84 57.84 73.53 36.76 77.45 72.72 79.41 75.00 81.86
vie_Latn 72.06 28.92 65.69 67.65 78.92 56.86 81.86 79.63 77.45 79.90 87.25
war_Latn 64.22 50.00 76.96 58.82 75.98 28.92 71.57 75.76 81.86 83.82 84.80
wol_Latn 45.59 42.16 55.88 29.90 50.98 22.06 48.53 51.03 50.00 47.55 57.84
xho_Latn 49.51 27.45 52.94 43.14 61.27 23.04 50.98 62.79 63.73 54.90 67.16
yid_Hebr 40.20 20.59 44.12 39.22 50.49 24.02 46.57 55.91 51.96 59.31 62.75
yor_Latn 45.10 31.86 46.57 29.41 45.10 23.04 44.12 49.22 49.02 57.35 58.82
yue_Hani 69.61 82.35 84.80 73.53 83.33 60.78 81.86 79.66 82.35 85.78 85.29
zsm_Latn 66.18 58.82 82.35 75.49 78.43 48.04 76.96 80.74 82.35 85.78 89.71
zul_Latn 55.88 24.51 49.51 50.98 66.18 28.43 51.96 69.02 65.69 69.61 74.02
Avg 57.14 42.29 63.57 54.21 66.85 32.15 61.88 69.51 69.90 70.04 75.67
Table 3: 3-shot accuracy with KNN or ICL (Part II). Methods with ‘(KNN)’ denote that KNN is used with retrieved samples, otherwise ICL is used.
Appendix B KNN Performance Across Layers
We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 2 and 3. As shown, layer 21 of MaLA500 and layer 11 of Glot500 achieve the best performance across layers. Therefore, the retrieved results based on these two layers are used in the baselines.
Figure 2: Results of 10-shot KNN (K-Nearest Neighbors) with Glot500 as retriever across layers.
Figure 3: Results of 10-shot KNN (K-Nearest Neighbors) with MaLA500 as retriever across layers.
Appendix C Effect of $k$
We conduct additional experiments to analyze the impact of the parameter $k$, with the results presented in Figure 4. Our findings indicate that XAMPLER performs optimally when $k=10$. However, as $k$ exceeds 10, there is a slight decrease in performance. This trend may be attributed to the possibility that increasing $k$ leads to fewer hard negatives for training the retrieval.
Figure 4: In-context learning with XAMPLER with different $k$.

### v2

@@
 Abstract
-Recent studies have shown that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving high-quality in-context examples, significantly improves in-context learning of English.
-However, adapting these methods to other languages, especially low-resource ones, presents challenges due to the scarcity of available cross-lingual retrievers and annotated data.
-In this paper, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data.
-XAMPLER first trains a retriever with positive/negative English samples, which are constructed based on the predictions of the multilingual large language model for in-context learning.
-Then, the trained retriever is directly employed to retrieve English examples as few-shot examples for in-context learning of target languages.
-Experiments on the massively multilingual text classification benchmark of SIB200 with 176 languages demonstrate that XAMPLER substantially improves the in-context learning performance across languages.
+Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning of English.
+However, adapting these methods to other languages, especially low-resource ones, poses challenges due to the scarcity of cross-lingual retrievers and annotated data.
+Thus, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data.
+XAMPLER first trains a retriever based on Glot500, a multilingual small language model, using positive and negative English examples constructed from the predictions of a multilingual large language model, i.e., MaLA500.
+Leveraging the cross-lingual capacity of the retriever, it can directly retrieve English examples as few-shot examples for in-context learning of target languages.
+Experiments on the multilingual text classification benchmark SIB200 with 176 languages show that XAMPLER substantially improves the in-context learning performance across languages.
 Our code is available at https://github.com/cisnlp/XAMPLER.
+XAMPLER: Learning to Retrieve Cross-Lingual In-Context Examples
+Peiqin Lin1,2, André F.
+T.
+Martins3,4,5, Hinrich Schütze1,2 1Center for Information and Language Processing, LMU Munich 2Munich Center for Machine Learning 3Instituto Superior Técnico, Universidade de Lisboa (Lisbon ELLIS Unit) 4Instituto de Telecomunicações  5Unbabel linpq@cis.lmu.de
 1 Introduction
@@
 The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals using labeled data.
+Figure 1: XAMPLER involves three steps: 1.
+Data Construction: given a query in English $q_{i}$, we divide the candidate examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2.
+Retriever Fine-tuning: we fine-tune the retriever based on Glot500 (Imani et al., 2023) using the constructed data; 3.
+In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples as few-shots for in-context learning.
+During training, XAMPLER is English-only, but for evaluation via in-context learning, it extends to any of the 500+ languages supported in MaLA500 and Glot500.
 Utilizing off-the-shelf retrievers has been further validated as an effective approach in multilingual settings (Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023).
 However, this method encounters limitations when applied to low-resource languages.
-Existing multilingual retrievers, e.g., (Reimers and Gurevych, 2020), cover a limited number of languages (i.e., 50+), and language-model-based retrievers (Hu et al., 2020) struggle to effectively align distant languages (Cao et al., 2020; Liu et al., 2023).
+Existing multilingual retrievers, e.g., SBERT (Reimers and Gurevych, 2020), cover a limited number of languages (i.e., 50+), and language-model-based retrievers (Hu et al., 2020) struggle to effectively align distant languages (Cao et al., 2020; Liu et al., 2023).
 Additionally, relying on off-the-shelf retrievers might lead to sub-optimal performance.
@@
 To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval).
-Specifically, given an English query and an English candidate training example as input, we employ in-context learning with MaLA500 (Lin et al., 2024b), a massively multilingual LLM covering 534 languages, to predict the label of the query.
-Based on the correctness of the prediction, we classify the candidate example as either positive or negative.
-Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a massively multilingual language model covering 534 languages, with the objective of minimizing the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023).
+As shown in Fig.
+1, given an English query $q_{i}$ and an English example from the candidate pool $D_{i}$, we employ in-context learning with MaLA500 (Lin et al., 2024b), a 10B multilingual LLM covering 534 languages, to predict the label of the query.
+Based on the correctness of the prediction, we classify the candidate example as either positive or negative, i.e., $D_{i}^{pos}$ and $D_{i}^{neg}$.
+Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a multilingual small language model covering 534 languages, aiming to minimize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023).
 Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language.
-The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for further in-context learning.
-Our extensive experiments across a range of 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning performance across languages.
+The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for in-context learning.
+Experiments across 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
 2 Approach
 2.1 Problem Definition
-Given an input query $q_{i}$ in any language, our aim is to retrieve tailored few-shot examples from the sample pool $D$ to enhance the in-context learning of $q_{i}$.
-Given the scarcity of annotated data in low-resource languages, we introduce XAMPLER, namely, Cross-Lingual Example Retrieval.
-On one hand, we leverage in-domain English examples as the sample pool $D$, from which we retrieve cross-lingual examples in English for $q_{i}$ in any target language.
-On the other hand, we only consider $q_{i}$ sourced from English training data to train the task-specific retriever, and then directly apply the retriever for evaluation across different languages.
+Given an input query $q_{i}$ in any language, our objective is to enhance in-context learning for predicting the label of $q_{i}$ by retrieving tailored few-shot examples from the pool of candidate examples $D$.
+Due to the scarcity of annotated data in low-resource languages, we introduce XAMPLER, namely, Cross-Lingual Example Retrieval.
+On one hand, we leverage in-domain English examples as the pool of candidate examples $D$, from which we retrieve cross-lingual examples in English for $q_{i}$ in any target language.
+On the other hand, we only consider $q_{i}$ sourced from English training data to train the task-specific retriever, which is then directly applied for evaluation across languages.
 2.2 Data Construction
-To train the task-specific retriever aimed at retrieving informative samples for the given query $q_{i}$, we consider contrastive learning, which requires both positive and negative examples for each query $q_{i}$.
-We define examples as positive when the LLM accurately predicts the ground truth of $q_{i}$ while utilizing the candidate as a one-shot example appended to $q_{i}$ for in-context learning.
+To train the task-specific retriever aimed at retrieving informative examples for the given query $q_{i}$, we consider contrastive learning, which requires both positive and negative examples for each query $q_{i}$.
+We define examples as positive when the LLM accurately predicts the ground truth of $q_{i}$ while utilizing the example as a one-shot example appended to $q_{i}$ for in-context learning.
 Conversely, examples are categorized as negative if the LLM’s prediction deviates from the ground truth.
@@
 (2022), we mitigate this by selecting the top $k$ similar examples as candidates.
-We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] This paper uses the distiluse-base-multilingual-cased-v1 version for SBERT.
+We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] We use version distiluse-base-multilingual-cased-v1.
 for candidate selection.
@@
 After obtaining the candidate-query pairs $\{(q_{i},d_{i,1}),\cdots,(q_{i},d_{i,k})\}$, we conduct 1-shot in-context learning with MaLA500 (Lin et al., 2024b) to predict the class of the $q_{i}$ given the candidate $d_{i,j}$, resulting in a predicted label $\hat{y}_{i,j}$.
-If MaLA500 correctly predicts the label of $q_{i}$ (i.e., $\hat{y_{i,j}}=y_{i}$), we consider the candidate $d_{i,j}$ as a positive sample ($d^{+}_{i,j}$); otherwise, it is treated as a negative sample ($d^{-}_{i,j}$).
-Finally, we divide $D_{i}$ into sets of positive and negative samples, denoted as $D_{i}^{pos}$ and $D_{i}^{neg}$, respectively.
+If MaLA500 correctly predicts the label of $q_{i}$ (i.e., $\hat{y}_{i,j}=y_{i}$), we consider the candidate $d_{i,j}$ as a positive example ($d^{+}_{i,j}$); otherwise a negative example ($d^{-}_{i,j}$).
+Finally, we divide $D_{i}$ into sets of positive and negative examples, denoted as $D_{i}^{pos}$ and $D_{i}^{neg}$, respectively.
 2.3 Retriever Fine-tuning
-We utilize a contrastive loss function (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between positive example pairs while minimizing the similarity between negative example pairs.
-We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant costs of fine-tuning a LLM.
+We utilize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between $q_{i}$ and $x_{i,j}$ if $x_{i,j}$ is a positive example while minimizing the similarity if $x_{i,j}$ is a negative example.
+We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant cost of fine-tuning an LLM.
+We train for 50 epochs with a learning rate of 2e-5.
+Due to the multilingual nature of Glot500, the fine-tuned retriever can be effectively transferred to retrieve in-context examples for other languages.
 2.4 In-Context Learning
-When employing in-context learning across languages, where $q_{i}$ can be in any language, we utilize the fine-tuned task-specific retriever to retrieve a few cross-lingual examples in English tailored to $q_{i}$.
-The retrieved examples are then appended to $q_{i}$ as input for MaLA500 (Lin et al., 2024b) to predict the label of $q_{i}$ through in-context learning.
+At test time, when employing in-context learning across languages, where $q_{i}$ can be in any language, we use the fine-tuned task-specific retriever to retrieve a few cross-lingual examples in English tailored to $q_{i}$.
+The retrieved examples are appended to $q_{i}$ as input for MaLA500 (Lin et al., 2024b) to predict the label of $q_{i}$ through in-context learning.
 3 Experiment
@@
 Benchmark
-We evaluate XAMPLER on a massively multilingual text classification benchmark, i.e., SIB200 (Adelani et al., 2023).
-SIB200 involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography.
-Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500.
-For a detailed language list, refer to §A.
+We evaluate XAMPLER on a massively multilingual text classification benchmark, SIB200 (Adelani et al., 2023).
+SIB200 involves seven classes: science/technology, travel, politics, sports, health, entertainment, and geography.
+Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §A).
+The English training set contains 701 samples, and each language has 204 samples for evaluation.
 Our evaluation framework follows the prompt template used in Lin et al.
@@
 We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources.
-MaLA500 takes in the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set.
+MaLA500 takes the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set.
 We measure the performance with accuracy.
 Baselines
-We compare XAMPLER with three types of baselines.
-  • Random Sampling. We randomly select examples from the training set $D$.
+Random Sampling. We randomly select examples from the English candidate pool $D$.
 Off-the-shelf Retriever.
@@
 Cross-lingual Transfer.
-We delve into another category of baselines called cross-lingual transfer, which also harnesses English data.
-In this approach, the language model is fine-tuned with English data and subsequently deployed for evaluation across target languages.
-Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines are denoted as Glot500 XLT and MaLA500 XLT.
+Cross-lingual transfer is another baseline exploiting English data.
+In this approach, the multilingual language model is fine-tuned with English data and then deployed for evaluation across target languages.
+Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines denoted as Glot500 XLT and MaLA500 XLT.
 For Glot500, we opt for full-parameter fine-tuning.
-For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the parameters of the LoRA modules with prompt tuning.
-In addition, we consider K-Nearest Neighbors (KNN) with XAMPLER and the baselines with retrievers for comparison.
-Specifically, we adopt majority voting based on the labels of the retrieved examples with the given retriever.
+For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the LoRA parameters with prompt tuning.
+K-Nearest Neighbors.
+We consider K-Nearest Neighbors (KNN) with the fine-tuned task-specific retriever of XAMPLER and the baselines with retrievers for comparison.
+Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
 KNN ICL
@@
 XAMPLER 67.05 75.97 70.04 73.59 79.80 75.67
-Table 1: Average macro-accuracy across all 176 languages, 117 languages in the Latin script and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines with K-Nearest Neighbors (KNN) or In-Context Learning (ICL).
-The provided results are all based on the 3-shot setting.
+Table 1: Average macro-accuracy across all 176 languages, 117 languages in Latin scripts and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines.
+KNN results are only provided for retriever-based methods.
+Results are based on the 3-shot setting.
 3.2 Main Results
@@
 For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever.
-Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%, respectively.
+Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%.
 Among all the methods, XAMPLER achieves the highest performance, surpassing the second-best method, MaLA500 XLT, by 5.77%.
@@
 3.4 Effect of In-Context Learning
-In Table 1, we compare the performance of XAMPLER with KNN against XAMPLER with ICL.
-Notably, XAMPLER with ICL exhibits a superior performance, outperforming KNN by 5.63%.
-Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages written in non-Latin scripts.
+In Table 1, we compare XAMPLER with KNN against that with ICL.
+Notably, XAMPLER with ICL surpasses XAMPLER with KNN by 5.63%.
+Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages in non-Latin.
 This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
-We further compare the performance of XAMPLER with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 1.
+We further compare XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2.
 Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%.
-Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL reaches its peak with only 2 examples.
+Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL achieves impressive results with only 2 examples.
 Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%.
 These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
-Figure 1: KNN (K-Nearest Neighbors) vs. ICL (In-Context Learning) with different number of shots. X-axis: the number of retrieved examples. Y-axis: Macro-average accuracy.
+Figure 2: KNN (K-Nearest Neighbors) vs. ICL (In-Context Learning) with different number of shots. X-axis: number of shots. Y-axis: Macro-average accuracy.
 4 Related Work
@@
 Hence, Rubin et al.
-(2022) delves into learning-based approaches: if an LLM finds an example useful, the retriever should be encouraged to retrieve it.
+(2022) delve into learning-based approaches: if an LLM finds an example useful, the retriever should be encouraged to retrieve it.
 This approach enables direct training of the retriever using signals derived from query and example pairs in the task of interest.
-Retrieval-based in-context learning has demonstrated effectiveness in natural language understanding (Liu et al., 2022; Ye et al., 2023), reasoning (Zhang et al., 2023), knowledge-based question answering (Liu et al., 2022; Ye et al., 2023), and text generation (Rubin et al., 2022).
-A study closely related to ours is conducted by Shi et al.
-(2022), which trains a cross-lingual exemplar retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task.
+Several works (Shi et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023) extend these methods to non-English languages.
+A study closely related to ours is Shi et al.
+(2022), which trains a cross-lingual example retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task.
 However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al.
@@
 In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language.
-By relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language.
+Relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language.
 Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.
-Limitation
-In this paper, we utilize Glot500 as the base model for training XAMPLER’s retriever, MaLA500 as the LLM for in-context learning, and SIB200 as the benchmark to assess the efficacy of XAMPLER across multiple languages.
-We opted not to consider alternative models and benchmarks due to the absence of massively multilingual models and benchmarks.
+Limitations
+We did not consider other models and benchmarks due to the absence / unavailability of massively multilingual ones.
 Additionally, while it is acknowledged that English may not universally serve as the optimal source language for cross-lingual transfer across all target languages (Lin et al., 2019; Wang et al., 2023; Lin et al., 2024a), our study does not explore the selection of different source languages due to the predominant availability of training data in English for many tasks.
@@
 Appendix B KNN Performance Across Layers
-We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 2 and 3.
+We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 3 and 4.
 As shown, layer 21 of MaLA500 and layer 11 of Glot500 achieve the best performance across layers.
@@
 However, as $k$ exceeds 10, there is a slight decrease in performance.
-This trend may be attributed to the possibility that increasing $k$ leads to fewer hard negatives for training the retrieval.
+This trend may be attributed to the possibility that increasing $k$ leads to fewer hard negatives for training the retriever.
 Figure 5: In-context learning with XAMPLER with different $k$.

### v3

@@
 Abstract
-Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning of English.
+Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning for English.
 However, adapting these methods to other languages, especially low-resource ones, poses challenges due to the scarcity of cross-lingual retrievers and annotated data.
@@
 Leveraging the cross-lingual capacity of the retriever, it can directly retrieve English examples as few-shot examples for in-context learning of target languages.
-Experiments on the multilingual text classification benchmark SIB200 with 176 languages show that XAMPLER substantially improves the in-context learning performance across languages.
+Experiments on two multilingual text classification benchmarks, namely SIB200 with 176 languages and MasakhaNEWS with 16 languages, demonstrate that XAMPLER substantially improves the in-context learning performance across languages.
 Our code is available at https://github.com/cisnlp/XAMPLER.
@@
 Rather than relying on random selection, these studies advocate for retrieving examples tailored to the input query, resulting in notable performance enhancements in in-context learning.
-The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals using labeled data.
+The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals (whether an example is helpful) using labeled data.
 Figure 1: XAMPLER involves three steps: 1.
-Data Construction: given a query in English $q_{i}$, we divide the candidate examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2.
+Data Construction: given a query in English $q_{i}$, we divide the candidate English examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2.
 Retriever Fine-tuning: we fine-tune the retriever based on Glot500 (Imani et al., 2023) using the constructed data; 3.
-In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples as few-shots for in-context learning.
-During training, XAMPLER is English-only, but for evaluation via in-context learning, it extends to any of the 500+ languages supported in MaLA500 and Glot500.
+In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples $D_{j}$ as few-shots for in-context learning.
+For training, XAMPLER requires English data only.
+Once trained, the model can be applied to any of the 500 languages covered by MaLA500/Glot500 without any need for (often unavailable) labeled low-resource data.
 Utilizing off-the-shelf retrievers has been further validated as an effective approach in multilingual settings (Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023).
@@
 The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for in-context learning.
-Experiments across 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
+Experiments across 176 languages on SIB200 and 16 languages on MasakhaNEWS show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
 2 Approach
@@
 for candidate selection.
-Based on our experiments detailed in Section C, we set $k=10$.
+Based on our experiments detailed in Section B, we set $k=10$.
 The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
@@
 We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant cost of fine-tuning an LLM.
-We train for 50 epochs with a learning rate of 2e-5.
+We train for 50 epochs using the AdamW optimizer with a learning rate of 2e-5 and a batch size of 16.
 Due to the multilingual nature of Glot500, the fine-tuned retriever can be effectively transferred to retrieve in-context examples for other languages.
@@
 Benchmark
-We evaluate XAMPLER on a massively multilingual text classification benchmark, SIB200 (Adelani et al., 2023).
-SIB200 involves seven classes: science/technology, travel, politics, sports, health, entertainment, and geography.
-Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §A).
-The English training set contains 701 samples, and each language has 204 samples for evaluation.
+We evaluate XAMPLER on two text classification benchmarks: SIB200 (Adelani et al., 2023a) and MasakhaNEWS (Adelani et al., 2023b).
+The partitioning for these datasets was predefined by the respective benchmarks.
+SIB200 is a massively multilingual text classification benchmark with seven classes.
+Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §C).
+The English training set contains 701 samples, with 204 evaluation samples per language.
+MasakhaNEWS is a news classification task for 16 African languages, covering six topics.
+The English training set contains 3.31k samples, with 175 to 948 evaluation samples per language.
 Our evaluation framework follows the prompt template used in Lin et al.
-(2024b): ‘The topic of the news [sent] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth.
+(2024b): ‘The topic of the news [sentence] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth.
 [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample.
@@
 Baselines
+We compare XAMPLER with the following retrieving strategies:
 Random Sampling. We randomly select examples from the English candidate pool $D$.
-Off-the-shelf Retriever.
-We utilize SBERT (Reimers and Gurevych, 2019), a model covering 50+ languages trained with parallel corpora based on mBERT (Devlin et al., 2019).
-Additionally, we employ two massively multilingual language models, namely Glot500 and MaLA500, as retrievers, denoted as Glot500 RET and MaLA500 RET, respectively.
+Multilingual Language Models.
+We use two massively multilingual language models, Glot500 and MaLA500, as retrievers.
 Tailored examples are retrieved based on the cosine similarity between the sentence representations of the candidate and the query.
@@
 For MaLA500, we adopt a position-weighted mean pooling method on the selected layer, assigning higher weights to later tokens (Muennighoff, 2022).
-We use K-Nearest Neighbors (KNN) to select the layer that performs best across layers (see §B).
+We use K-Nearest Neighbors to select the layer that performs best across layers (see §A).
 The selected layers for Glot500 and MaLA500 are 11 and 21, respectively.
-Cross-lingual Transfer.
-Cross-lingual transfer is another baseline exploiting English data.
-In this approach, the multilingual language model is fine-tuned with English data and then deployed for evaluation across target languages.
-Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines denoted as Glot500 XLT and MaLA500 XLT.
-For Glot500, we opt for full-parameter fine-tuning.
-For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the LoRA parameters with prompt tuning.
-K-Nearest Neighbors.
-We consider K-Nearest Neighbors (KNN) with the fine-tuned task-specific retriever of XAMPLER and the baselines with retrievers for comparison.
-Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
-KNN ICL
-Latin Non-Latin Avg Latin Non-Latin Avg
-Random - - - 56.38 58.66 57.14
-SBERT 44.90 37.13 42.29 63.90 62.93 63.57
-Glot500 RET 51.16 60.26 54.21 65.21 70.09 66.85
-MaLA500 RET 31.90 32.63 32.15 61.02 63.58 61.88
-Glot500 XLT - - - 67.09 74.30 69.51
-MaLA500 XLT - - - 69.15 71.39 69.90
-XAMPLER 67.05 75.97 70.04 73.59 79.80 75.67
-Table 1: Average macro-accuracy across all 176 languages, 117 languages in Latin scripts and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines.
-KNN results are only provided for retriever-based methods.
-Results are based on the 3-shot setting.
+Off-the-shelf Retriever.
+We also employ three off-the-shelf retrievers trained on parallel data: SBERT (Reimers and Gurevych, 2019), LaBSE (Feng et al., 2022), and Multilingual E5 (Wang et al., 2024).
+[^2] https://huggingface.co/intfloat/multilingual-e5-large
+We set the number of shots as the number of classes and evaluate under two settings: the label-aware setting, where one shot is provided per class, and the label-agnostic setting, where the most similar examples are retrieved regardless of their labels.
+SIB200 MasakhaNEWS
+Label-Aware Label-Agnostic Label-Aware Label-Agnostic
+Random 65.24 61.68 72.32 72.39
+Glot500 66.60 68.55 73.35 73.01
+MaLA500 66.75 66.25 73.39 71.58
+SBERT 67.13 66.59 73.24 72.8
+LaBSE 68.51 73.69 72.54 73.29
+Multilingual E5 69.09 74.61 73.63 72.61
+XAMPLER 70.18 75.91 75.02 73.85
+Table 1: Average macro-accuracy across the evaluated languages on SIB200 and MasakhaNEWS using XAMPLER compared to the baselines.
 3.2 Main Results
@@
 Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning.
-Leveraging an off-the-shelf retriever notably boosts performance.
-For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever.
-Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%.
-Among all the methods, XAMPLER achieves the highest performance, surpassing the second-best method, MaLA500 XLT, by 5.77%.
-This highlights the effectiveness of training a task-specific retriever solely with English data.
-Moreover, XAMPLER outperforms MaLA500 XLT by 4.44% on languages written in Latin scripts and by 8.41% on languages written in non-Latin scripts.
-It shows that XAMPLER can provide greater benefits for languages with non-Latin scripts, which are relatively isolated from English written in Latin.
-3.3 Effect of Task-Specific Retriever
-The results from various retrievers utilizing KNN, as shown in Table 1, indicate that XAMPLER’s fine-tuned retriever excels at retrieving more examples within the same classes as the query in the target language.
-Notably, XAMPLER with KNN outperforms the second-best retriever, Glot500 RET, by a notable margin, i.e., 15.83%.
-This superiority enables XAMPLER to leverage majority label bias in in-context learning (Zhao et al., 2021), thereby enhancing overall performance.
-3.4 Effect of In-Context Learning
-In Table 1, we compare XAMPLER with KNN against that with ICL.
-Notably, XAMPLER with ICL surpasses XAMPLER with KNN by 5.63%.
-Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages in non-Latin.
-This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
-We further compare XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2.
+Multilingual language models, which retrieve examples based on semantic similarity learned during pretraining, slightly outperform the random baseline.
+Leveraging an off-the-shelf retriever further improves performance, with Multilingual E5 emerging as the top performer among them.
+Among all the methods, XAMPLER achieves the highest performance in in-context learning.
+Specifically, on SIB200, XAMPLER surpasses Multilingual E5 by 1.09% in the label-aware setting and 1.30% in the label-agnostic setting, and by 1.93% and 1.24% on MasakhaNEWS, respectively.
+Our two established practices to reduce resource requirements, i.e., selecting the top-k similar examples as candidates using existing off-the-shelf retrievers and using a smaller base model for retriever training, ensures that XAMPLER operates efficiently.
+For example, on the SIB200 benchmark (701 samples, 10 candidates per sample) using an NVIDIA GeForce GTX 1080 Ti (11GB), retriever fine-tuning took just 1.5 hours, and retrieving similar examples added only 0.1 seconds per query during inference.
+3.3 Ablation Study
+To further assess the effectiveness of XAMPLER, we conduct the following ablation studies:
+Method Accuracy (%)
+XLT (Glot500) 69.51
+XLT (MaLA500) 69.90
+MT 74.50
+KNN 72.85
+XAMPLER 75.91
+Table 2: Performance comparison of XAMPLER with ablation methods on the SIB200 benchmark.
+Cross-lingual Transfer (XLT).
+Cross-lingual transfer is another approach that utilizes English data.
+In this method, the multilingual language model is fine-tuned on English data and then evaluated across target languages.
+Both Glot500 and MaLA500 are included, with their respective cross-lingual transfer methods denoted as XLT (Glot500) and XLT (MaLA500).
+For Glot500, we perform full-parameter fine-tuning with a batch size of 16 and a learning rate of 1e-5.
+For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we update only the LoRA parameters with prompt tuning.
+The learning rate is set to 1e-3, weight decay is 0.1, the maximum sequence length is 128, and the batch size is 16.
+The optimizer used is AdamW.
+Machine Translation (MT).
+We translate the examples retrieved by XAMPLER from English to the target language and use these translated examples as few-shot examples for in-context learning.
+For translation, we use the distilled 600M variant of NLLB-200 (Costa-jussà et al., 2022).
+K-Nearest Neighbors (KNN).
+We consider KNN with the fine-tuned task-specific retriever of XAMPLER for comparison.
+Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
+The results of the ablation studies are shown in Table 2.
+XAMPLER outperforms XLT (Glot500) and XLT (MaLA500) by 6.40% and 6.01%, respectively, demonstrating that in-context learning, when combined with informative cross-lingual few-shot examples, is superior to traditional cross-lingual transfer methods via fine-tuning.
+Translating English examples into target languages results in slightly worse results than XAMPLER by 1.41%.
+It demonstrates that XAMPLER benefits more from English in-context examples than in-language examples.
+A comparison between XAMPLER and KNN shows that XAMPLER performs better by 3.06%.
+We also compared XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2.
 Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%.
@@
 This approach enables direct training of the retriever using signals derived from query and example pairs in the task of interest.
-Several works (Shi et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023) extend these methods to non-English languages.
+Several works (Winata et al., 2021; Shi et al., 2022; Winata et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023; Cahyawijaya et al., 2024) extend these methods to non-English languages.
 A study closely related to ours is Shi et al.
@@
 Relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language.
-Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.
+Experiments on SIB200 and MasakhaNEWS show that XAMPLER outperforms previous methods by a notable margin.
 Limitations
@@
 Additionally, while it is acknowledged that English may not universally serve as the optimal source language for cross-lingual transfer across all target languages (Lin et al., 2019; Wang et al., 2023; Lin et al., 2024a), our study does not explore the selection of different source languages due to the predominant availability of training data in English for many tasks.
-Appendix A Detailed Results
-The language list of SIB200 and the results of XAMPLER and the compared baselines are shown in Table 2 and Table 3.
-Random SBERT (KNN) SBERT Glot500 RET (KNN) Glot500 RET MaLA500 RET (KNN) MaLA500 RET Glot500 XLT MaLA500 XLT XAMPLER (KNN) XAMPLER
-ace_Latn 61.76 40.20 64.71 50.00 69.12 28.92 69.12 65.12 66.18 75.49 75.98
-acm_Arab 55.39 79.90 79.41 64.22 73.04 35.29 62.75 76.81 74.51 76.47 79.90
-afr_Latn 65.20 75.00 83.33 60.78 78.92 43.63 72.06 76.10 78.92 82.84 86.27
-ajp_Arab 58.82 76.47 79.41 62.25 71.57 32.84 63.24 75.44 72.55 82.84 83.33
-als_Latn 64.71 46.57 75.49 63.73 73.53 36.76 67.65 76.76 81.86 65.20 83.33
-amh_Ethi 53.92 29.41 58.33 46.57 55.39 22.55 57.35 68.46 61.27 71.57 73.53
-apc_Arab 59.80 79.41 79.41 62.25 74.51 31.37 62.25 77.52 75.00 82.84 84.80
-arb_Arab 58.33 83.33 81.37 61.76 70.10 36.76 67.16 77.40 78.43 78.43 85.78
-ary_Arab 56.37 79.41 78.92 55.39 67.65 25.98 56.37 74.90 67.16 80.39 80.88
-arz_Arab 55.39 74.51 74.02 62.25 69.61 33.33 63.73 77.08 71.08 78.92 80.88
-asm_Beng 61.27 22.55 61.76 59.80 77.45 22.55 66.67 74.93 76.96 78.92 84.80
-ast_Latn 70.59 73.04 81.86 63.73 79.90 49.51 81.37 80.81 84.31 75.49 89.71
-ayr_Latn 37.25 24.51 40.69 32.84 41.67 19.12 42.65 49.93 48.53 38.24 50.98
-azb_Arab 44.61 26.47 47.55 41.18 59.31 23.04 47.55 65.56 58.82 72.55 73.04
-azj_Latn 66.18 58.33 78.92 72.55 76.47 20.10 68.63 78.73 87.25 75.49 85.78
-bak_Cyrl 58.33 50.98 74.02 59.80 72.55 28.92 63.73 77.87 75.00 72.55 80.88
-bam_Latn 36.76 31.37 42.65 31.86 43.14 19.61 41.18 49.71 44.12 46.08 46.08
-ban_Latn 66.18 44.61 70.59 53.92 70.59 31.86 72.06 73.01 72.06 75.98 82.35
-bel_Cyrl 64.22 38.24 72.06 59.31 76.96 43.63 73.04 77.16 79.90 69.12 84.31
-bem_Latn 42.16 28.43 43.63 32.35 50.00 18.14 39.71 56.89 58.33 60.78 61.76
-ben_Beng 63.73 20.10 61.27 60.29 75.00 32.35 69.12 73.95 80.88 79.90 82.35
-bjn_Latn 63.73 31.37 62.75 59.31 70.10 33.33 70.59 69.80 69.61 76.96 81.86
-bod_Tibt 44.61 09.31 33.33 38.24 48.53 19.61 47.55 62.79 47.06 54.90 59.31
-bos_Latn 70.10 55.39 77.45 72.55 79.90 46.08 75.00 81.99 83.33 71.57 87.75
-bul_Cyrl 65.69 75.00 82.84 67.16 80.39 52.45 78.92 78.87 81.86 79.90 85.29
-cat_Latn 69.61 74.51 84.31 64.22 74.51 61.27 80.88 77.97 86.76 71.57 88.73
-ceb_Latn 66.18 45.59 73.53 63.73 80.88 31.37 72.55 78.80 81.86 82.35 85.29
-ces_Latn 69.61 53.92 76.96 63.24 76.47 55.88 78.92 77.25 86.27 76.96 88.24
-cjk_Latn 37.75 28.43 42.16 26.96 45.10 21.57 44.61 48.41 53.43 44.61 47.06
-ckb_Arab 57.84 17.65 61.27 54.90 72.06 23.53 61.76 74.46 75.00 79.90 81.37
-cmn_Hani 68.63 85.29 83.82 75.98 80.88 62.75 78.92 78.55 82.84 87.25 88.73
-crh_Latn 58.82 67.65 75.00 54.41 70.59 14.71 58.33 67.84 71.08 64.71 75.98
-cym_Latn 63.24 24.02 65.69 49.02 73.04 27.94 70.10 72.50 77.45 72.55 81.37
-dan_Latn 68.14 47.06 78.92 67.65 79.90 56.37 78.43 79.68 84.80 86.27 89.71
-deu_Latn 70.59 83.82 86.27 70.10 77.45 57.84 80.88 78.48 87.25 80.88 87.75
-dyu_Latn 41.67 28.92 45.59 27.94 50.49 24.51 46.57 45.17 46.57 43.14 48.53
-dzo_Tibt 31.37 09.31 20.59 47.55 41.18 16.67 32.84 62.21 33.33 68.63 52.94
-ell_Grek 67.65 31.37 76.47 59.80 72.55 40.20 72.06 74.71 79.90 75.98 83.82
-eng_Latn 70.59 84.31 87.25 79.41 86.27 69.61 81.86 83.65 85.29 89.22 91.18
-epo_Latn 65.69 57.84 75.98 61.27 76.96 33.33 68.63 75.20 77.94 78.92 82.84
-est_Latn 61.76 44.61 66.67 65.69 75.49 30.88 66.18 74.07 76.47 67.65 81.37
-eus_Latn 57.35 47.55 71.08 66.67 76.47 29.90 68.14 76.08 70.10 66.67 82.84
-ewe_Latn 38.24 26.47 39.22 27.94 39.71 20.59 38.73 47.65 46.57 50.98 53.92
-fao_Latn 56.86 31.86 58.33 45.10 63.24 29.90 60.78 76.15 69.61 82.35 81.86
-fij_Latn 41.18 32.35 49.02 31.86 46.08 21.08 42.65 55.74 51.47 50.00 55.39
-fin_Latn 65.20 36.76 71.57 62.75 74.02 48.53 74.51 75.98 80.88 76.96 83.82
-fon_Latn 39.22 20.10 41.67 33.33 45.59 20.10 42.65 44.41 45.59 41.67 48.53
-fra_Latn 69.61 86.27 84.31 60.78 79.41 63.24 79.90 81.20 85.29 85.78 89.22
-ful_Latn 37.75 35.29 48.04 23.53 48.04 21.57 45.59 45.54 45.10 47.06 50.49
-fur_Latn 60.29 63.73 72.55 54.90 67.16 36.76 70.59 67.75 73.53 75.49 79.90
-gla_Latn 54.90 17.16 52.45 39.22 59.80 25.00 61.76 59.00 60.29 59.31 63.73
-gle_Latn 57.84 23.53 57.84 49.02 65.20 31.86 62.75 62.94 69.12 70.10 74.02
-glg_Latn 72.06 76.96 81.86 67.65 80.88 53.43 79.90 79.36 84.31 81.37 87.75
-grn_Latn 53.92 62.75 69.61 43.14 61.76 27.45 60.78 69.85 64.71 76.47 75.00
-guj_Gujr 59.80 13.73 53.43 68.14 73.04 32.84 65.20 77.87 73.53 81.86 86.27
-hat_Latn 61.76 50.49 72.06 59.31 74.02 25.49 68.14 73.55 81.86 79.41 82.84
-hau_Latn 56.86 23.04 55.88 47.06 60.78 23.53 57.35 62.33 70.10 58.82 70.10
-heb_Hebr 45.59 26.96 47.06 58.33 58.33 29.90 43.63 72.60 56.86 75.00 75.00
-hin_Deva 60.29 12.75 56.86 67.65 71.08 34.80 65.20 76.84 76.47 84.31 85.78
-hne_Deva 56.86 15.69 54.41 55.88 70.10 28.43 66.67 70.07 75.49 79.41 80.88
-hrv_Latn 70.10 52.94 78.43 73.53 82.35 51.47 74.51 81.86 85.78 71.57 88.24
-hun_Latn 64.71 38.24 66.18 69.61 81.86 50.49 72.06 81.20 84.80 83.82 85.78
-hye_Armn 67.16 16.18 67.16 60.78 70.59 31.37 68.14 76.62 77.45 80.88 83.33
-ibo_Latn 57.35 37.25 63.24 49.02 70.10 23.53 56.86 69.09 72.06 74.02 79.41
-ilo_Latn 61.76 50.49 69.12 47.55 68.14 32.35 65.69 70.81 77.45 75.00 79.90
-ind_Latn 69.61 64.22 82.84 75.98 81.37 50.98 75.98 80.34 83.82 88.24 89.71
-isl_Latn 60.78 31.86 67.65 55.39 69.61 24.51 62.75 73.04 75.49 77.94 78.92
-ita_Latn 71.08 82.84 88.24 68.14 81.37 64.22 77.94 78.80 86.76 79.90 88.24
-jav_Latn 64.22 40.69 71.08 56.37 75.00 30.88 70.10 73.19 75.00 79.90 82.35
-jpn_Jpan 71.08 76.96 87.25 68.14 76.96 65.69 82.84 78.33 80.88 83.33 87.75
-kab_Latn 28.43 20.10 27.94 27.45 36.27 22.06 29.41 38.70 27.94 31.37 37.75
-kac_Latn 35.29 28.92 41.18 40.20 42.65 20.10 36.76 54.09 45.10 40.69 46.57
-kam_Latn 38.73 30.88 47.55 34.31 46.08 17.65 34.31 46.99 50.49 46.08 50.00
-kan_Knda 58.82 15.69 56.37 64.22 71.57 24.02 59.80 74.73 73.04 75.00 77.94
-kat_Geor 65.20 15.69 58.82 63.73 75.98 27.94 70.10 77.79 75.98 72.55 83.33
-kaz_Cyrl 61.76 43.14 68.63 69.12 74.51 31.37 64.71 75.47 78.43 67.16 79.41
-kbp_Latn 37.25 25.00 43.14 34.31 44.61 17.65 44.12 49.83 47.06 41.18 48.04
-kea_Latn 64.71 61.27 78.92 48.53 75.49 33.82 75.98 66.81 80.39 74.51 79.41
-khm_Khmr 69.12 34.31 69.61 58.82 76.47 38.24 74.51 76.42 79.41 82.84 86.76
-kik_Latn 46.57 33.33 51.47 43.63 50.98 19.12 44.12 55.47 52.94 58.82 59.31
-kin_Latn 45.59 26.96 50.00 39.22 51.47 19.61 48.53 57.01 67.65 64.71 64.22
-kir_Cyrl 58.82 42.65 67.65 68.63 75.98 24.51 55.39 75.93 75.00 63.73 77.45
-kmb_Latn 36.76 24.51 41.67 28.43 42.65 24.02 42.65 44.80 48.04 46.57 50.98
-kmr_Latn 51.96 27.45 58.82 50.00 69.12 20.10 56.37 63.55 66.18 52.94 70.59
-kon_Latn 49.02 50.49 63.24 44.61 63.73 20.10 54.41 58.85 61.27 57.84 64.22
-kor_Hang 69.61 81.37 85.29 62.25 78.43 48.53 78.43 77.33 79.90 82.35 87.25
-lao_Laoo 64.71 33.33 69.12 58.33 74.02 30.39 70.59 77.33 75.00 79.90 84.31
-lij_Latn 64.22 61.76 77.45 50.49 69.61 35.29 66.67 70.07 76.96 70.59 80.39
-lim_Latn 61.76 68.63 79.41 55.88 72.55 38.24 72.55 69.49 65.20 71.08 75.98
-lin_Latn 48.04 46.57 61.27 48.53 61.27 20.59 54.90 64.95 58.33 63.24 68.63
-Table 2: 3-shot accuracy with KNN or ICL (Part I). Methods with ‘(KNN)’ denote that KNN is used with retrieved samples, otherwise ICL is used.
-Random SBERT (KNN) SBERT Glot500 RET (KNN) Glot500 RET MaLA500 RET (KNN) MaLA500 RET Glot500 XLT MaLA500 XLT XAMPLER (KNN) XAMPLER
-lit_Latn 61.27 41.18 64.71 62.25 72.55 29.41 63.24 78.04 79.41 69.61 85.29
-lmo_Latn 63.24 57.35 73.04 48.53 66.67 33.33 66.18 66.25 73.53 72.55 77.45
-ltz_Latn 62.75 60.78 76.96 56.37 72.55 35.29 70.10 69.26 80.39 72.06 79.41
-lua_Latn 38.24 42.65 50.49 36.27 45.10 24.51 47.06 55.51 52.45 47.55 53.43
-lug_Latn 38.24 29.41 41.67 33.82 50.00 20.10 40.69 53.28 51.47 52.45 58.33
-luo_Latn 37.75 26.47 40.69 31.86 43.63 18.63 44.12 51.03 50.00 55.39 59.31
-lus_Latn 48.53 41.18 57.84 39.71 50.98 30.39 51.96 60.83 61.76 65.69 68.14
-lvs_Latn 64.22 39.22 68.63 62.25 73.53 31.86 64.22 78.87 79.90 71.08 82.84
-mai_Deva 57.35 17.16 50.49 62.25 69.61 28.92 59.31 76.91 68.63 83.82 78.92
-mal_Mlym 57.84 15.69 52.45 58.33 67.16 26.47 59.80 73.33 68.63 77.94 78.92
-mar_Deva 59.31 15.69 55.88 60.78 75.00 27.94 63.73 75.96 77.94 73.04 79.41
-min_Latn 67.65 44.61 66.67 58.82 72.55 32.84 74.02 70.42 78.43 78.43 79.90
-mkd_Cyrl 68.63 57.84 77.94 67.16 75.98 49.02 75.49 76.50 80.88 72.55 83.82
-mlt_Latn 68.14 50.00 72.55 60.29 77.45 31.86 67.65 77.87 81.37 78.92 84.31
-mon_Cyrl 58.82 21.08 59.31 63.73 73.53 25.00 62.25 75.12 71.57 79.41 82.84
-mos_Latn 37.75 39.22 46.08 23.04 39.22 21.08 42.65 47.06 45.10 44.61 50.49
-mri_Latn 52.94 22.55 54.41 18.63 46.08 17.16 46.08 54.49 60.78 59.80 68.63
-mya_Mymr 53.92 16.67 49.02 53.92 66.18 23.04 50.98 73.68 54.90 74.02 80.39
-nld_Latn 68.63 82.84 85.29 75.00 81.37 56.37 81.37 78.24 86.27 87.25 88.73
-nno_Latn 68.14 46.57 73.53 64.22 75.49 42.65 74.51 78.63 83.82 86.27 89.22
-npi_Deva 63.24 15.20 59.31 67.16 74.51 29.41 70.10 77.99 75.49 82.35 87.75
-nso_Latn 41.67 30.39 52.45 39.22 50.49 21.57 43.63 60.71 55.39 54.41 58.82
-nya_Latn 45.10 35.29 52.94 50.49 61.76 23.04 49.02 68.36 58.82 66.67 68.14
-oci_Latn 67.65 69.61 81.86 58.33 70.10 44.12 77.45 76.50 84.31 74.02 85.78
-orm_Latn 34.80 13.73 34.80 40.69 42.16 15.69 35.29 50.59 46.08 42.65 48.53
-ory_Orya 51.47 29.41 47.06 59.31 67.65 23.53 57.35 73.55 62.25 80.88 78.92
-pag_Latn 55.88 63.24 72.06 59.31 71.57 26.47 62.25 73.31 75.98 80.39 80.88
-pan_Guru 57.35 14.22 53.92 51.96 67.16 26.47 60.78 70.44 73.04 71.08 75.98
-pap_Latn 65.20 63.73 75.49 58.82 76.96 34.31 70.10 72.23 78.43 77.94 80.39
-pes_Arab 66.67 37.25 67.65 69.61 76.96 34.31 71.08 77.72 79.41 87.75 86.76
-plt_Latn 52.94 29.41 57.35 42.65 60.29 21.08 50.49 64.22 69.61 48.04 67.16
-pol_Latn 68.14 78.43 87.25 67.16 78.92 57.84 78.43 77.97 84.80 75.00 85.29
-por_Latn 72.06 81.86 87.75 60.78 81.37 59.31 83.33 79.71 85.78 85.78 90.69
-prs_Arab 65.20 32.84 64.22 62.75 76.96 31.86 71.57 78.70 78.92 86.76 86.27
-pus_Arab 49.51 25.98 52.94 53.43 59.80 23.53 50.98 69.34 62.25 60.78 71.08
-quy_Latn 49.02 29.41 52.45 42.65 52.94 19.61 53.92 59.31 57.84 56.37 62.25
-ron_Latn 67.65 58.33 77.94 66.18 75.98 55.39 76.96 78.14 83.33 80.88 86.76
-run_Latn 43.14 22.06 44.61 41.18 49.51 18.14 45.59 51.86 60.78 62.75 66.18
-rus_Cyrl 69.61 83.82 86.27 72.06 79.90 56.86 77.45 79.53 81.37 86.27 89.71
-sag_Latn 43.14 39.71 49.02 39.71 54.41 20.59 45.59 52.50 49.02 56.37 59.31
-san_Deva 51.96 15.20 52.94 48.04 59.80 23.53 59.31 70.34 60.29 66.67 68.63
-scn_Latn 68.14 55.39 77.45 49.51 71.57 34.31 73.53 67.50 74.51 72.06 80.39
-sin_Sinh 61.27 22.06 62.25 66.18 73.53 26.47 65.20 74.29 72.55 79.90 82.84
-slk_Latn 66.67 49.51 70.59 63.73 74.02 45.10 75.98 77.99 81.86 74.02 84.31
-slv_Latn 63.73 49.02 74.51 64.22 74.02 51.96 75.00 75.51 78.92 66.67 81.86
-smo_Latn 54.41 32.84 63.24 36.76 62.75 18.63 52.94 69.14 63.73 66.18 73.04
-sna_Latn 45.10 24.51 47.06 42.16 51.47 21.57 40.20 59.31 60.29 55.39 62.75
-snd_Arab 44.12 29.90 49.51 50.49 52.45 20.10 44.61 65.83 47.55 59.80 67.16
-som_Latn 42.65 20.59 42.65 49.02 59.80 17.16 42.16 56.52 63.24 50.00 58.82
-sot_Latn 45.10 24.51 49.02 43.63 58.33 21.57 48.04 63.63 57.84 64.71 67.65
-spa_Latn 71.57 84.80 85.78 67.16 78.92 59.80 80.39 79.14 85.29 82.84 90.20
-srd_Latn 63.73 64.22 75.98 49.51 70.59 34.80 71.08 66.10 66.67 61.76 78.43
-srp_Cyrl 68.63 55.88 79.41 65.69 82.84 47.55 79.41 78.53 82.35 67.65 80.88
-ssw_Latn 45.10 24.51 44.12 40.69 56.86 26.96 51.96 63.19 59.80 59.80 63.24
-sun_Latn 68.14 43.14 70.10 63.24 78.43 29.90 73.53 75.93 81.37 82.84 85.29
-swe_Latn 66.18 46.08 72.55 71.57 77.94 59.31 77.94 79.73 85.29 84.80 87.75
-swh_Latn 61.27 28.92 60.29 56.86 66.67 12.75 54.90 72.94 76.96 64.22 75.98
-szl_Latn 62.75 67.65 75.98 54.41 70.10 40.69 71.57 68.65 75.49 63.73 77.45
-tam_Taml 54.90 15.20 50.98 64.22 71.08 27.45 62.25 75.37 75.98 80.88 78.92
-tat_Cyrl 61.76 46.57 71.57 63.73 75.00 24.51 68.63 79.17 78.92 70.59 82.35
-tel_Telu 57.84 18.63 50.98 65.69 70.59 25.98 60.29 74.53 74.51 84.80 81.37
-tgk_Cyrl 60.29 26.96 61.27 63.73 73.53 35.29 63.73 77.08 80.88 82.35 80.39
-tgl_Latn 68.14 38.73 69.61 63.24 79.41 30.88 70.10 78.21 83.82 83.33 86.76
-tha_Thai 65.20 25.98 56.86 65.20 75.49 41.18 67.65 77.43 70.10 79.90 86.76
-tir_Ethi 44.12 29.41 51.47 42.16 48.53 18.63 47.06 55.91 50.00 55.39 57.35
-tpi_Latn 65.20 53.92 75.98 57.84 70.59 27.94 64.71 76.94 74.51 81.37 84.80
-tsn_Latn 49.02 31.37 49.02 35.78 50.49 20.10 42.16 56.76 56.86 54.90 59.80
-tso_Latn 45.59 25.98 44.12 38.73 51.47 22.06 50.98 56.23 55.39 55.88 62.25
-tuk_Latn 53.43 50.00 68.63 51.47 69.61 14.22 50.98 73.33 66.67 61.27 77.94
-tum_Latn 42.16 26.96 47.06 32.35 56.86 24.02 43.63 62.55 52.94 65.20 64.71
-tur_Latn 68.14 82.84 86.27 68.63 76.96 15.69 56.86 78.06 81.37 73.53 84.80
-uig_Arab 45.10 12.75 40.20 55.39 53.43 18.63 37.25 72.18 49.51 72.55 69.12
-ukr_Cyrl 66.18 59.31 77.45 67.16 79.90 56.86 79.41 78.63 81.86 76.47 86.27
-umb_Latn 33.82 29.90 37.75 33.82 46.57 22.06 39.71 48.97 49.02 47.06 49.51
-urd_Arab 55.88 29.90 54.90 59.80 69.12 36.27 66.67 74.29 69.12 64.22 75.98
-uzb_Latn 57.35 38.24 62.25 59.80 73.04 24.02 62.75 73.14 76.96 60.78 77.45
-vec_Latn 67.65 68.14 82.84 57.84 73.53 36.76 77.45 72.72 79.41 75.00 81.86
-vie_Latn 72.06 28.92 65.69 67.65 78.92 56.86 81.86 79.63 77.45 79.90 87.25
-war_Latn 64.22 50.00 76.96 58.82 75.98 28.92 71.57 75.76 81.86 83.82 84.80
-wol_Latn 45.59 42.16 55.88 29.90 50.98 22.06 48.53 51.03 50.00 47.55 57.84
-xho_Latn 49.51 27.45 52.94 43.14 61.27 23.04 50.98 62.79 63.73 54.90 67.16
-yid_Hebr 40.20 20.59 44.12 39.22 50.49 24.02 46.57 55.91 51.96 59.31 62.75
-yor_Latn 45.10 31.86 46.57 29.41 45.10 23.04 44.12 49.22 49.02 57.35 58.82
-yue_Hani 69.61 82.35 84.80 73.53 83.33 60.78 81.86 79.66 82.35 85.78 85.29
-zsm_Latn 66.18 58.82 82.35 75.49 78.43 48.04 76.96 80.74 82.35 85.78 89.71
-zul_Latn 55.88 24.51 49.51 50.98 66.18 28.43 51.96 69.02 65.69 69.61 74.02
-Avg 57.14 42.29 63.57 54.21 66.85 32.15 61.88 69.51 69.90 70.04 75.67
-Table 3: 3-shot accuracy with KNN or ICL (Part II). Methods with ‘(KNN)’ denote that KNN is used with retrieved samples, otherwise ICL is used.
-Appendix B KNN Performance Across Layers
+Acknowledgements
+This work was funded by DFG (SCHU 2246/14-1), the European Research Council (DECOLLAGE, ERC-2022-CoG #101088763), EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research, and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
+Appendix A KNN Performance Across Layers
 We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 3 and 4.
@@
 Figure 4: Results of 10-shot KNN (K-Nearest Neighbors) with MaLA500 as retriever across layers.
-Appendix C Effect of $k$
+Appendix B Effect of $k$
 We conduct additional experiments to analyze the impact of the parameter $k$, with the results presented in Figure 5.
@@
 Figure 5: In-context learning with XAMPLER with different $k$.
+Appendix C Detailed Results
+The language list of SIB200 and the results of XAMPLER and the compared baselines are shown in Table 3 and Table 4.
+The language list of MasakhaNEWS and the results of XAMPLER and the compared baselines are shown in Table 5.
+Label-Aware Label-Agnostic
+Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER
+ace_Latn 72.55 75.00 72.06 74.51 72.55 74.51 72.06 63.73 71.57 70.10 70.10 76.47 80.39 76.96
+acm_Arab 62.25 64.71 65.20 70.10 66.18 68.63 75.00 61.76 71.08 65.69 78.92 75.98 75.98 79.90
+afr_Latn 77.45 76.47 77.94 79.90 80.39 81.86 79.41 70.59 79.41 78.43 83.82 88.24 81.86 87.25
+ajp_Arab 64.71 67.65 67.16 68.14 69.61 68.14 74.51 64.71 72.06 64.71 77.94 76.47 77.94 84.31
+als_Latn 74.51 74.51 77.94 78.92 77.45 79.41 75.00 68.63 75.98 75.00 80.88 83.82 82.84 84.31
+amh_Ethi 56.37 59.31 60.29 60.29 61.76 60.78 68.63 57.35 61.27 60.78 60.29 67.16 67.16 71.08
+apc_Arab 63.73 67.65 65.20 68.63 68.14 71.08 75.00 61.76 72.06 62.25 78.43 79.90 76.47 85.78
+arb_Arab 65.69 69.61 68.63 69.61 72.55 72.06 77.45 65.20 74.02 70.10 76.96 81.37 75.98 83.82
+ary_Arab 58.82 65.20 63.24 65.69 64.71 66.18 74.51 57.84 71.57 60.78 71.57 71.57 73.53 80.39
+arz_Arab 63.24 66.18 64.71 66.18 67.65 65.20 73.53 62.25 69.61 61.76 76.47 78.92 76.96 82.84
+asm_Beng 70.59 72.06 73.04 72.06 75.98 75.49 73.53 69.12 76.47 72.55 64.22 78.92 76.96 83.82
+ast_Latn 79.90 76.47 81.37 79.90 79.41 80.39 82.84 74.02 78.43 81.37 85.78 84.31 82.84 89.71
+ayr_Latn 39.71 45.59 43.14 46.08 44.61 45.59 42.16 44.12 47.55 47.06 46.08 46.57 53.43 52.45
+azb_Arab 47.55 50.49 50.98 50.98 52.45 54.41 64.71 48.04 61.27 49.51 46.57 60.29 61.76 70.59
+azj_Latn 77.45 76.47 77.45 77.94 78.92 82.35 77.94 71.57 79.41 76.47 78.92 82.84 83.82 84.80
+bak_Cyrl 69.12 75.49 71.57 72.55 72.06 74.51 75.00 66.18 73.53 69.61 77.94 75.00 78.92 80.88
+bam_Latn 44.12 46.08 48.04 47.06 49.51 50.49 52.94 43.14 47.55 44.61 46.08 50.00 57.35 49.02
+ban_Latn 75.00 74.51 76.47 77.94 78.92 78.43 79.90 69.61 73.53 78.43 75.98 83.33 84.31 82.35
+bel_Cyrl 76.96 76.96 75.00 78.43 77.94 78.92 76.47 71.08 77.94 74.02 76.47 84.80 81.86 85.78
+bem_Latn 50.49 52.94 52.45 49.51 52.45 54.41 53.43 47.55 58.82 50.98 50.49 54.41 65.69 62.25
+ben_Beng 72.06 72.06 67.65 69.61 73.53 72.55 79.41 65.20 75.49 68.14 66.18 77.94 77.94 81.86
+bjn_Latn 71.57 73.04 72.55 72.06 73.53 73.53 76.47 71.08 73.04 74.51 70.10 80.39 79.90 83.33
+bod_Tibt 53.92 50.49 49.51 50.98 56.86 50.98 50.98 48.53 51.96 51.96 36.76 56.37 50.98 59.31
+bos_Latn 78.43 78.43 75.98 78.92 81.37 81.86 79.90 72.06 82.35 77.94 80.39 82.84 82.84 89.22
+bul_Cyrl 77.94 78.92 77.45 79.41 78.92 77.94 78.92 72.06 78.92 78.43 81.86 85.78 80.88 85.78
+cat_Latn 78.43 77.45 80.88 83.82 80.39 83.33 81.86 74.02 81.86 79.41 84.31 86.76 84.31 88.24
+ceb_Latn 75.98 75.98 76.96 76.47 78.92 76.47 79.90 71.08 75.98 76.96 74.51 84.80 81.37 86.27
+ces_Latn 75.49 76.96 77.94 76.96 78.43 79.41 77.94 69.61 77.45 76.96 78.92 85.78 82.35 88.73
+cjk_Latn 44.12 44.12 44.12 44.12 47.06 46.57 43.14 40.69 46.08 49.02 44.61 50.00 55.88 47.06
+ckb_Arab 69.12 71.08 66.67 69.61 69.12 72.06 75.00 63.24 73.53 67.16 66.18 63.24 75.98 81.86
+cmn_Hani 76.96 77.45 81.86 79.90 79.41 79.90 86.27 69.12 81.37 77.94 82.84 79.90 80.88 89.71
+crh_Latn 71.08 69.12 67.65 72.06 74.02 72.55 72.55 62.25 74.02 69.61 73.04 75.98 75.49 75.00
+cym_Latn 77.45 76.96 75.98 75.98 80.39 78.43 77.94 71.57 78.43 79.90 70.59 85.78 84.31 78.92
+dan_Latn 81.37 82.84 80.88 82.35 79.90 83.82 84.80 75.49 82.35 82.84 82.35 86.76 82.84 89.71
+deu_Latn 80.39 80.88 83.82 83.82 82.35 83.33 83.33 74.51 78.43 84.31 85.78 86.27 85.29 86.76
+dyu_Latn 51.96 51.96 50.98 50.98 51.47 56.37 49.02 47.55 49.02 49.51 48.04 54.41 57.84 50.49
+dzo_Tibt 31.86 39.22 35.78 33.33 40.69 37.75 50.00 34.31 43.14 37.25 22.06 48.53 34.80 54.90
+ell_Grek 74.02 75.00 78.92 77.45 77.45 79.41 76.47 72.06 76.47 74.02 75.98 78.92 80.39 83.82
+eng_Latn 82.84 84.31 85.29 85.78 84.31 84.80 86.27 73.04 84.80 85.78 86.27 85.29 87.75 91.67
+epo_Latn 73.53 76.96 75.00 74.51 76.96 77.45 79.90 71.08 78.43 77.94 76.96 86.27 84.80 82.84
+est_Latn 68.63 73.04 71.08 71.57 74.02 75.49 75.00 67.16 75.98 71.57 72.55 79.90 80.39 79.90
+eus_Latn 59.80 70.10 69.12 69.61 69.61 73.04 75.49 63.73 75.49 68.63 69.61 79.90 86.27 83.82
+ewe_Latn 42.65 45.59 44.61 47.55 48.04 47.06 48.04 41.67 42.16 43.14 47.55 49.51 54.90 53.43
+fao_Latn 62.75 66.18 68.14 64.71 69.61 69.61 73.53 59.31 70.10 63.24 65.69 77.94 77.45 83.82
+fij_Latn 43.14 42.65 46.57 45.10 45.59 48.04 50.49 44.12 49.51 47.55 47.55 54.41 60.29 53.43
+fin_Latn 75.49 75.00 75.49 74.51 75.98 77.94 77.94 67.16 78.92 74.51 74.51 82.84 80.39 83.33
+fon_Latn 43.14 46.57 46.57 46.57 50.98 52.45 42.16 41.18 48.53 42.65 46.08 50.49 57.35 48.04
+fra_Latn 78.92 78.43 78.43 80.88 81.37 81.86 83.82 72.06 79.41 81.86 80.39 86.76 82.84 89.22
+ful_Latn 45.59 47.06 47.55 50.49 52.94 50.98 43.63 44.61 46.08 50.00 51.96 57.35 55.88 52.94
+fur_Latn 69.61 68.63 71.08 69.61 70.59 73.04 75.49 63.73 68.14 70.59 75.98 74.51 80.88 79.41
+gla_Latn 63.24 57.84 64.71 65.20 68.63 68.14 62.75 60.29 63.24 68.63 59.80 75.49 74.51 67.16
+gle_Latn 65.20 68.14 71.08 71.08 71.08 72.55 66.18 66.18 72.55 67.65 64.71 80.88 81.37 74.02
+glg_Latn 79.41 80.88 79.41 79.90 77.94 81.37 84.80 71.57 83.33 81.86 86.76 84.80 87.75 86.76
+grn_Latn 63.73 64.22 64.22 67.16 65.69 69.12 70.10 61.76 67.16 66.18 69.61 71.08 75.00 77.94
+guj_Gujr 73.53 73.04 70.10 69.61 77.45 73.04 79.90 69.61 73.04 67.65 62.25 77.45 78.92 85.29
+hat_Latn 70.10 72.55 75.00 73.53 75.00 76.96 76.47 65.20 77.45 73.53 73.04 82.84 81.37 82.35
+hau_Latn 68.14 65.69 67.65 63.73 67.16 68.14 69.12 58.82 68.63 66.67 61.27 76.47 74.02 69.12
+heb_Hebr 47.55 50.00 45.10 49.02 52.45 54.41 62.75 47.55 61.76 45.59 50.00 65.69 68.63 77.45
+hin_Deva 68.63 71.08 69.61 68.63 73.53 76.47 78.92 69.12 78.43 70.59 63.24 80.39 79.90 85.29
+hne_Deva 67.16 74.02 69.12 68.63 75.00 77.45 74.51 63.24 75.00 70.59 62.75 81.37 80.39 80.88
+hrv_Latn 79.41 80.88 78.92 77.94 78.92 82.35 80.39 73.04 80.88 76.47 80.39 84.80 82.35 86.76
+hun_Latn 76.47 75.00 77.45 77.45 77.45 76.96 80.88 71.08 76.96 76.96 77.94 86.27 78.43 87.25
+hye_Armn 74.02 75.49 72.06 74.02 74.51 75.98 76.47 66.67 74.51 71.57 70.59 79.90 80.39 84.80
+ibo_Latn 71.08 73.53 72.06 73.53 73.53 74.02 75.00 66.18 71.57 69.61 71.57 79.90 83.33 80.88
+ilo_Latn 66.67 69.12 71.08 69.61 73.53 75.49 74.51 62.75 71.57 71.57 74.51 76.47 83.33 78.92
+ind_Latn 79.41 81.86 80.88 80.88 81.37 82.35 84.80 74.51 80.88 83.33 83.33 84.80 85.78 90.69
+isl_Latn 69.12 69.12 69.61 70.59 71.57 73.04 74.02 65.20 68.63 67.65 69.61 78.92 72.55 81.37
+ita_Latn 80.88 79.90 82.35 82.84 83.33 83.33 84.80 73.53 82.35 83.33 86.27 87.25 86.27 90.69
+jav_Latn 72.06 74.02 72.55 74.51 75.00 75.49 77.45 69.61 75.98 74.02 76.47 77.45 78.43 85.78
+jpn_Jpan 78.43 78.43 80.88 80.88 83.82 81.86 83.33 74.51 77.94 81.86 79.90 86.76 80.88 86.27
+kab_Latn 34.31 36.27 37.75 37.75 38.73 39.71 34.31 33.33 38.24 33.33 33.82 40.69 40.20 35.78
+kac_Latn 40.20 41.67 41.18 39.71 45.10 44.12 42.65 39.71 40.69 42.65 47.06 47.06 54.90 46.08
+kam_Latn 41.18 43.63 40.69 44.12 47.55 47.06 44.12 40.69 50.98 44.61 44.61 51.47 55.39 49.51
+kan_Knda 69.61 72.06 66.18 69.61 71.08 73.53 74.02 65.20 74.51 68.63 64.71 76.96 77.45 77.45
+kat_Geor 76.47 75.49 77.45 74.51 77.45 77.45 78.92 74.02 79.41 76.96 64.71 83.82 78.92 83.82
+kaz_Cyrl 72.55 73.04 73.04 73.53 75.00 75.49 77.45 64.22 75.49 70.59 73.53 80.88 79.41 82.35
+kbp_Latn 43.14 44.12 48.04 48.04 47.55 48.53 49.51 43.63 42.16 46.57 47.55 44.61 50.98 47.55
+kea_Latn 74.51 75.98 75.00 77.94 75.00 76.96 79.41 68.14 77.45 75.49 82.84 81.37 82.35 78.92
+khm_Khmr 76.96 76.47 75.49 74.51 76.47 77.94 80.88 71.57 77.45 74.51 75.00 78.92 79.41 85.29
+kik_Latn 50.00 53.92 55.39 52.94 54.90 55.39 58.82 48.04 54.90 52.94 53.43 57.35 64.71 58.82
+kin_Latn 49.51 51.96 51.96 50.49 51.47 51.96 56.37 49.02 55.39 50.49 48.04 64.71 72.06 62.75
+kir_Cyrl 68.14 69.61 69.61 70.59 72.06 74.51 72.06 64.22 73.04 64.71 69.61 75.98 77.94 78.43
+kmb_Latn 40.20 41.67 42.16 43.14 44.61 46.08 43.63 43.63 47.06 44.12 43.14 48.53 53.43 50.98
+kmr_Latn 58.33 61.76 61.76 61.27 65.20 66.18 67.16 55.88 69.61 61.27 57.84 74.02 76.96 69.61
+kon_Latn 58.33 64.71 61.27 63.24 65.69 67.16 63.24 56.37 61.76 59.80 61.76 59.31 74.51 64.22
+kor_Hang 76.96 78.92 80.88 83.82 81.37 81.86 81.37 75.98 80.39 81.86 87.25 86.27 85.29 88.73
+lao_Laoo 72.55 70.59 74.51 72.06 73.04 70.59 75.98 66.67 73.53 74.02 70.59 76.47 78.43 84.31
+lij_Latn 73.04 73.04 72.06 74.02 74.51 75.49 75.98 67.16 72.55 73.53 78.43 76.96 78.92 80.39
+lim_Latn 77.45 76.47 75.98 79.41 77.94 77.45 75.98 68.14 78.92 76.96 81.86 83.82 82.35 78.43
+lin_Latn 59.31 57.84 60.78 63.24 60.29 59.31 62.75 52.45 63.24 60.78 65.20 61.27 67.65 66.67
+Table 3: 7-shot accuracy with on SIB200.
+Label-Aware Label-Agnostic
+Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER
+lit_Latn 66.67 68.63 67.65 69.12 71.57 68.63 74.02 66.18 72.55 70.10 70.10 78.43 80.39 86.27
+lmo_Latn 70.10 72.06 73.04 72.06 72.55 73.53 75.98 66.18 75.00 73.53 74.02 79.41 78.92 77.45
+ltz_Latn 74.02 72.55 74.51 73.53 77.94 75.49 76.47 69.61 78.92 78.43 76.47 83.82 83.82 80.88
+lua_Latn 48.53 48.53 50.00 49.02 51.96 50.98 49.51 48.04 50.49 50.00 54.41 57.35 62.25 53.92
+lug_Latn 43.14 47.55 49.51 45.10 47.06 48.04 46.57 44.61 49.51 46.57 45.59 53.43 61.27 56.37
+luo_Latn 45.10 46.08 47.55 48.04 49.51 51.96 48.04 44.61 43.63 47.55 49.51 54.90 56.37 58.82
+lus_Latn 55.39 54.41 56.86 56.37 59.31 59.31 58.33 51.47 53.43 57.84 57.84 64.71 63.73 69.12
+lvs_Latn 67.16 67.65 72.06 70.10 73.04 72.55 75.98 63.73 74.51 70.59 75.98 79.41 80.88 79.41
+mai_Deva 62.25 67.16 68.63 66.18 72.06 72.55 72.55 58.82 72.55 61.27 60.78 79.41 76.47 83.33
+mal_Mlym 65.69 66.18 66.18 63.73 68.14 69.12 75.49 62.25 69.61 65.20 54.90 74.02 75.00 80.39
+mar_Deva 69.12 72.06 71.57 72.06 74.51 74.02 75.49 66.18 72.55 70.10 67.65 79.41 76.96 80.88
+min_Latn 76.47 75.98 75.00 76.96 76.47 78.43 75.49 69.12 75.00 75.49 75.49 81.86 78.92 81.37
+mkd_Cyrl 74.51 75.49 77.45 76.47 76.47 78.43 77.94 70.10 78.92 77.94 77.45 82.35 81.37 82.35
+mlt_Latn 77.45 75.98 79.41 78.92 77.45 81.86 79.90 71.57 76.96 79.41 78.92 81.37 84.80 86.76
+mon_Cyrl 71.57 70.59 71.08 72.55 74.02 73.04 78.92 65.69 75.00 68.63 64.22 76.96 80.39 82.84
+mos_Latn 45.59 45.10 47.55 46.08 46.57 44.61 43.14 41.67 44.61 45.59 47.55 50.98 50.49 51.47
+mri_Latn 62.25 61.27 62.25 63.24 67.65 66.18 62.25 59.31 56.37 64.22 58.82 72.55 74.51 70.59
+mya_Mymr 63.73 59.31 61.27 64.22 63.24 67.65 71.08 59.80 59.80 54.90 59.80 70.10 67.16 76.96
+nld_Latn 80.39 81.37 81.37 85.29 84.31 85.29 84.80 74.51 83.82 83.82 84.80 85.78 86.76 89.22
+nno_Latn 76.47 76.96 77.94 80.39 79.41 79.41 83.33 74.02 75.49 78.43 81.86 85.78 82.35 90.69
+npi_Deva 72.55 73.04 70.59 71.57 75.98 75.49 77.45 69.61 78.43 71.57 63.73 82.35 81.37 87.25
+nso_Latn 50.98 53.43 51.47 56.37 55.88 57.35 55.88 48.53 56.37 50.49 53.92 59.80 64.71 59.80
+nya_Latn 54.90 55.39 58.33 53.92 60.29 57.84 64.22 52.45 62.25 55.88 54.90 69.61 71.08 69.61
+oci_Latn 79.41 77.94 78.43 80.39 77.94 80.39 79.41 68.63 76.47 78.43 80.88 83.33 80.88 85.29
+orm_Latn 38.24 38.24 40.20 37.75 41.18 40.69 36.76 37.25 39.22 39.71 37.75 43.14 55.39 46.08
+ory_Orya 62.75 65.20 60.78 57.84 60.78 64.71 72.55 59.31 67.65 58.33 58.33 69.61 68.14 80.88
+pag_Latn 67.16 68.63 68.14 71.08 72.06 70.10 75.00 60.29 71.08 68.63 75.98 72.55 76.96 80.39
+pan_Guru 61.27 67.16 63.73 64.22 65.20 65.20 72.06 59.80 67.65 62.25 58.82 73.04 74.51 78.43
+pap_Latn 74.02 75.49 76.47 75.49 75.98 78.43 78.43 70.59 75.49 76.96 78.43 79.90 84.31 80.39
+pes_Arab 75.00 75.98 76.47 75.00 77.45 75.98 83.33 73.53 77.45 75.49 73.53 82.35 82.35 87.75
+plt_Latn 57.84 65.20 60.29 58.82 62.25 63.73 55.88 58.82 64.71 60.29 58.33 72.55 75.00 63.73
+pol_Latn 77.94 80.88 78.43 81.37 78.92 79.90 81.86 74.51 79.41 80.39 83.82 82.35 81.37 87.25
+por_Latn 77.94 81.37 81.37 83.82 82.35 82.84 85.29 75.49 83.33 81.86 85.78 88.24 86.27 91.18
+prs_Arab 74.02 75.49 73.04 73.53 75.00 75.49 82.84 68.63 80.88 74.02 69.12 81.37 84.31 87.25
+pus_Arab 54.90 54.90 57.35 58.33 59.31 59.31 66.18 53.92 60.29 55.88 53.92 66.67 68.63 71.57
+quy_Latn 53.43 56.37 56.86 57.35 57.35 56.37 63.73 53.43 55.88 56.37 60.29 61.27 65.69 62.75
+ron_Latn 76.47 75.98 78.92 76.96 76.47 77.45 81.86 71.08 75.49 79.41 84.31 80.88 81.37 84.31
+run_Latn 48.04 49.02 52.94 47.55 52.94 53.92 54.41 47.06 48.53 49.51 50.98 67.16 70.10 65.69
+rus_Cyrl 79.41 79.90 78.92 83.82 81.37 81.86 82.35 70.59 80.88 79.90 85.78 85.78 85.29 90.20
+sag_Latn 52.94 50.00 53.43 52.45 53.43 52.94 57.84 49.51 53.92 49.51 50.49 51.47 54.90 59.31
+san_Deva 57.35 61.27 64.71 59.80 64.71 62.75 66.18 59.31 62.75 59.31 56.86 69.12 69.61 71.08
+scn_Latn 75.98 77.45 78.43 78.92 80.88 81.86 79.41 70.59 73.04 77.94 79.90 80.39 81.37 80.88
+sin_Sinh 69.61 72.55 66.18 70.10 72.06 72.55 73.53 66.18 70.59 69.12 67.65 77.94 75.49 81.86
+slk_Latn 73.53 75.00 75.49 75.49 78.43 77.94 78.92 69.12 79.90 75.49 76.47 83.82 83.33 84.31
+slv_Latn 74.02 75.49 75.00 74.02 77.45 77.45 75.98 68.63 73.53 76.47 74.51 83.33 80.88 84.80
+smo_Latn 61.27 62.25 67.16 65.20 68.14 69.61 66.67 60.78 65.69 64.71 64.22 73.04 71.57 73.53
+sna_Latn 54.41 54.90 55.39 52.94 55.39 55.39 50.00 50.00 54.41 50.98 51.47 60.29 70.10 63.24
+snd_Arab 48.04 50.00 50.00 53.43 54.41 57.84 56.37 50.49 59.31 50.00 53.92 63.24 65.69 66.67
+som_Latn 49.02 50.49 51.47 52.45 53.43 53.43 49.02 46.08 61.27 49.51 46.57 67.65 66.18 57.84
+sot_Latn 58.33 62.25 61.76 62.75 65.20 62.25 59.80 54.41 62.25 58.82 50.00 64.71 68.63 69.61
+spa_Latn 78.92 79.90 78.43 82.84 80.39 81.37 84.80 73.04 77.94 79.41 83.33 85.29 83.33 89.71
+srd_Latn 75.49 71.57 75.98 73.53 75.00 72.06 74.02 69.12 74.51 77.94 78.43 78.43 80.88 77.45
+srp_Cyrl 76.47 80.39 77.94 78.43 81.37 77.45 77.45 72.55 81.86 79.41 81.37 83.33 83.33 81.86
+ssw_Latn 50.98 52.45 56.37 53.43 55.88 59.31 60.29 53.43 58.82 55.39 49.51 65.20 69.12 63.73
+sun_Latn 77.94 75.49 75.98 79.41 80.39 79.41 80.88 74.02 77.45 76.96 78.92 83.33 80.88 86.76
+swe_Latn 75.49 76.47 76.47 79.41 79.90 79.41 83.82 71.08 76.47 78.92 75.98 81.86 82.84 86.76
+swh_Latn 65.20 62.75 64.71 63.24 66.18 65.20 68.63 61.27 65.69 60.78 59.80 73.53 72.06 77.45
+szl_Latn 72.55 73.04 75.49 74.02 75.98 75.49 76.96 67.65 71.08 71.57 75.49 78.92 81.86 75.00
+tam_Taml 68.14 67.16 68.63 68.14 67.65 69.12 74.02 62.75 67.16 68.14 57.35 75.98 75.98 78.92
+tat_Cyrl 72.06 75.00 72.55 72.55 74.51 75.98 76.47 67.65 75.98 73.53 70.59 82.84 77.45 81.37
+tel_Telu 66.18 69.61 66.18 64.22 72.06 70.10 75.00 59.31 73.04 64.22 65.69 75.98 75.98 81.86
+tgk_Cyrl 69.61 70.10 70.10 71.08 74.02 70.10 73.53 63.24 70.59 68.63 68.14 79.90 75.98 81.37
+tgl_Latn 76.47 79.41 79.41 78.92 78.92 79.90 81.37 71.57 78.92 80.39 74.02 84.80 84.80 87.75
+tha_Thai 76.96 78.43 75.49 76.47 76.47 78.43 78.43 70.59 78.43 72.06 65.69 79.41 79.90 85.78
+tir_Ethi 47.55 48.53 50.98 51.96 46.08 48.04 51.96 47.55 52.45 46.57 47.06 53.43 58.82 57.35
+tpi_Latn 76.47 78.43 79.90 79.41 79.90 80.39 85.78 72.55 77.45 75.00 78.92 80.88 82.35 85.29
+tsn_Latn 54.90 52.45 56.86 55.88 58.82 57.84 55.88 50.49 52.94 53.92 52.94 64.71 66.67 61.27
+tso_Latn 54.90 55.39 56.37 57.35 54.90 57.35 57.84 50.49 56.37 54.41 50.00 55.88 68.63 63.24
+tuk_Latn 61.76 66.18 65.20 67.16 70.10 66.18 69.61 60.78 72.55 61.27 66.18 75.49 75.98 75.98
+tum_Latn 50.00 53.43 53.43 50.98 52.94 55.88 53.43 50.00 55.88 49.02 49.51 62.75 70.59 67.16
+tur_Latn 77.45 75.98 76.47 78.92 75.98 78.43 81.86 67.65 79.41 74.51 84.80 83.33 79.90 85.29
+uig_Arab 44.12 49.02 43.63 45.59 50.00 53.43 62.25 45.10 50.98 42.16 42.65 58.33 53.43 66.67
+ukr_Cyrl 75.00 76.47 73.53 76.96 79.41 78.92 78.92 71.08 80.39 78.43 80.39 83.33 82.35 86.27
+umb_Latn 33.33 41.67 39.22 38.73 39.71 42.65 40.20 35.78 42.65 39.22 37.25 46.57 56.37 50.49
+urd_Arab 66.67 69.12 68.14 64.71 66.67 66.67 74.02 58.33 68.63 65.69 60.29 77.94 72.55 75.49
+uzb_Latn 69.61 70.59 68.14 69.61 71.57 71.57 72.55 65.20 75.49 67.16 65.69 81.37 77.45 77.94
+vec_Latn 77.45 76.96 78.43 79.41 80.88 80.39 78.43 73.53 76.47 78.43 81.37 81.86 80.88 81.37
+vie_Latn 81.37 79.90 81.86 79.90 82.84 82.84 84.80 74.51 79.41 82.35 69.61 85.78 85.78 87.25
+war_Latn 74.51 77.45 76.47 76.96 82.35 79.90 83.33 69.61 78.92 77.94 80.88 82.35 84.80 83.82
+wol_Latn 53.92 53.43 54.41 55.39 56.86 58.33 54.90 49.51 49.51 51.47 57.84 58.82 64.71 58.33
+xho_Latn 57.84 61.76 61.76 59.80 61.27 65.20 63.24 58.82 62.75 62.25 55.39 68.63 71.57 66.18
+yid_Hebr 47.06 48.04 49.51 53.43 49.02 54.41 51.47 47.06 50.00 49.02 45.59 57.35 60.78 62.75
+yor_Latn 47.55 50.49 51.47 49.51 52.94 57.35 50.00 45.59 48.53 50.00 51.47 62.25 62.75 57.84
+yue_Hani 78.43 81.37 83.82 82.84 83.82 86.27 84.31 71.08 81.86 80.88 83.82 85.29 81.86 85.29
+zsm_Latn 76.96 79.41 78.43 80.39 81.37 79.41 81.37 72.55 81.37 79.41 80.88 85.78 82.84 89.71
+zul_Latn 61.27 64.22 61.76 65.69 66.67 66.18 70.59 55.88 67.65 59.31 57.35 74.02 72.55 76.96
+Avg 65.24 66.60 66.75 67.13 68.51 69.09 70.18 61.68 68.55 66.25 66.59 73.69 74.61 75.91
+Table 4: 7-shot accuracy with on SIB200.
+Label-Aware Label-Agnostic
+Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER Random Glot500 MaLA500 SBERT LaBSE Multilingual E5 XAMPLER
+amh 76.06 76.33 78.46 78.19 77.13 77.66 81.38 78.99 78.72 72.34 81.65 75.00 77.39 79.52
+eng 76.05 82.17 81.96 81.12 81.22 81.12 82.81 75.32 81.43 81.54 81.86 83.02 83.02 86.50
+fra 85.31 84.60 85.07 83.18 83.18 83.18 85.31 84.12 81.75 83.89 80.57 81.99 79.62 80.81
+hau 63.94 63.20 62.45 63.01 63.20 64.31 65.43 61.15 61.52 59.48 62.08 63.94 64.87 66.91
+ibo 82.40 85.87 83.73 85.33 81.60 85.33 83.73 83.73 84.80 85.87 85.60 82.67 81.87 78.13
+lin 82.29 86.29 81.14 84.57 84.00 81.71 81.14 82.86 80.57 77.71 82.29 86.29 81.14 83.43
+lug 70.10 67.16 67.65 66.18 65.20 67.16 66.18 67.16 66.67 64.71 66.18 65.20 64.71 69.12
+orm 59.08 60.62 61.54 59.69 59.69 62.46 67.69 60.92 61.54 60.00 61.54 59.08 58.15 68.62
+pcm 93.44 94.10 94.43 94.10 93.44 95.08 95.41 93.11 94.10 93.11 93.77 93.44 93.11 93.11
+run 78.83 80.78 79.15 80.13 79.48 80.78 79.48 79.48 80.46 78.18 79.15 80.78 80.78 80.13
+sna 69.11 65.58 66.67 65.85 62.87 63.96 62.06 63.69 63.96 61.25 63.96 61.79 65.31 59.08
+som 62.01 62.37 64.52 63.08 62.37 63.80 63.80 60.57 62.37 63.44 61.29 64.87 63.08 58.42
+swa 74.34 72.90 71.94 72.42 73.38 72.66 76.02 70.26 72.90 70.26 72.18 75.30 73.14 75.30
+tir 51.84 55.88 56.25 58.82 56.25 57.72 65.44 57.35 61.03 54.41 61.76 61.40 59.56 65.07
+xho 48.15 50.51 54.21 52.53 53.20 55.89 57.91 52.86 50.51 55.22 46.80 54.55 53.54 55.56
+yor 84.15 85.30 85.01 83.57 84.44 85.30 86.46 86.74 85.88 83.86 84.15 83.29 82.42 81.84
+Avg 72.32 73.35 73.39 73.24 72.54 73.63 75.02 72.39 73.01 71.58 72.80 73.29 72.61 73.85
+Table 5: 6-shot accuracy with on MasakhaNEWS.